### v2

# Scalable Private Search with Wally

Abstract
This paper presents Wally  a private search system that supports efficient semantic and keyword search queries against large databases. When sufficient clients are making the queries, Wally performance is significantly better than previous systems. In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry. As a result, performance degraded proportionally with the number of entries in the database.
In Wally we get rid of this limitation. Specifically, for each query the server performs cryptographic operations only against a few database entries. We achieve these results by requiring each client to add a few fake queries, and sends each query via an anonymous network to the server at independently chosen random instants. Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake, Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
The number of fake queries each client makes depends inversely on the number of clients making queries. Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases. Concretely, Wally can serve $8$M requests at a rate of 3,000 queries per second. That is around 60x higher than the state-of-the-art scheme.
1 Introduction
Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server. This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo. These search engines improve the quality of our lives by promptly providing helpful information. Nevertheless, they require learning the client’s query to respond accurately to each query, hence providing no query privacy. In some situations, learning the query could reveal sensitive and personal information about the client. For example, consider the queries “High blood pressure with late-stage HIV?”, “Closest diabetes clinic from Albuquerque”, and “Best divorce lawyers ”. Revealing these queries to search engines could be damaging to the client. However, even if the queries do not appear sensitive, they may allow the server to learn personal information like medical information, marital status, sexuality, etc. Therefore, enabling query privacy is critical for search engines.
Private search engines opt to provide query privacy. However, these systems have significantly higher computation and communication overhead than the state-of-the-art insecure solutions. For example, in Tiptoe [HDCZ23], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries. The high computation is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in. Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick. We detail this trade off in a later section.. Additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
In this paper, we present Wally, a privacy preserving search engine that efficiently scales to performance constraints of large-scale search engines. Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword. To provide privacy, our key observation is that in search engines, there are always enough clients making queries. We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries. In our protocol, the server computation and the communication get amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy. In Wally the server learns noised differentially private histogram of anonymized queries from all clients. We emphasize that Tiptoe, is fully oblivious, i.e., the server learns nothing besides that a particular client is making a query. Therefore, the privacy guarantee of Tiptoe is stronger than Wally  however differential privacy is an accepted standard for strong privacy [DR+14]. Specifically, in Wally every client is guaranteed that all the statistical inferences are insensitive to their data. In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
At a high level, Wally works in epochs. Within each epoch, a fixed number of users make queries. The epoch size is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch. The server database is divided into $C$ clusters. Each query consists of a relevant cluster index, and the server response is all the entries in the cluster. The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries. The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
Even though queries are anonymized, they do not provide query privacy. The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (if queries arrive earlier than the others). To prevent this leakage, Wally implements the following two changes at the client.
To ensure that the server does not learn information by observing the access patterns, in Wally each client also makes fake queries. For each fake query, the client picks a randomly selected bucket. This ensures that the server cannot differentiate between real and fake queries. The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private. In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch. Note that increasing the number of clients decreases the average number of fake queries. The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases. We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
Wally also ensures that arrival time and order of the queries leak no information about the client to the server. To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch. Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
Anonymizing the queries, adding fake queries that are indistinguishable from real ones, and randomizing query order and time guarantees differential privacy.
As mentioned above, for each query, the server response contains all the entries in the cluster. For large clusters, this would result in very high response overhead. In Wally  we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead. Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE. The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores. This reduces the response because the size of encrypted scores is significantly smaller than the entries. For a keyword search, the client includes encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword. To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data. We propose various optimizations that significantly improve performance for both kinds of searches.
Our contributions are as follows:
  • We describe Wally’s design, which enables privacy-preserving large-scale search engines. Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt their queries.
  • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee. Generally, SHE computations have a high performance overhead. To achieve desired performance constraints, we propose several optimizations to SHE computation (distance computation and PIR) that could be of independent interest.
  • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
Limitation.
Wally requires decent number of clients always present in an epoch to make queries. We argue that this assumption aligns with current search engines, for example a hundred of millions of people make queries in Google search daily [Fle24]. We remark that for systems with few users, overhead due to fake queries could get quite high, therefore for these systems fully oblivious protocols like Tiptoe will be suitable.
2 Background
Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one. Matrices and vectors are represented as capital boldface letters and capital lowercase, respectively, e.g., $\mathbf{M}$ and $\mathbf{v}$. The entry-wise, Hadamard, product of two vectors is denoted as $\mathbf{v}\odot\mathbf{v}^{\prime}$. All logarithms are the natural log unless noted otherwise. We denote a random variable being sampled from a probability distribution $D$ as $a\leftarrow D$. All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set. We denote the integers as $\mathbb{Z}$ and the integers modulo a positive integer $Q$ as $\mathbb{Z}_{Q}$. Let $\mathcal{NB}(r,p)$ denote the negative binomial distribution parameterized by $r,p$. Its probability mass function is $Pr(X=k)=\binom{k+r-1}{k}(1-p)^{k}p^{r}$, $X\sim\mathcal{NB}(r,p)$, for every non-negative integer $k$. An important fact about the negative binomial distribution is that it is infinitely divisible, Definition 1.
Definition 1.
A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ iid variables. That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
2.1 Homomorphic encryption
We use the BFV [Bra12, FV12] SHE scheme [^2] BFV is also a fully homomorphic scheme, which can homomorphically compute any function, with large enough parameter sets. We only focus on its simpler SHE parameter regime in our paper. where each plaintext is a polynomial in $\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and $t$ is a prime or prime power. Wally can also be implemented with the BGV [BGV14] SHE scheme with similar performance. Both schemes are post-quantum SHE schemes based on the ring-learning-with-errors (RLWE) problem [Reg05, LPR10]. A BFV ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})\in R_{Q}^{2}$ reduced modulo $Q$ and $X^{n}+1$ for a power of two $n$, $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$. The number $n$ is a ring dimension, $n\in\{2^{12},2^{13}\}$ in our use cases, and $Q$ is a ciphertext modulus with $\log_{2}Q=83$ or $\log_{2}Q=155$ since we use 128-bit secure post-quantum security parameters [ADPS16].
A BFV ciphertext satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=\lfloor Q/t\rceil m+e$ where $s$ is a secret key polynomial with small coefficients, in $\{0,\pm 1\}$, $e$ is small error polynomial, $t$ is the plaintext modulus, and $m\in\mathbb{Z}_{t}[X]/(X^{n}+1)$ is the plaintext polynomial. This error polynomial’s entries grow as we compute more homomorphic operations on the ciphertext and we can only decrypt if $\|e\|_{\infty}<Q/2t$. We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
SHE parameters.
We choose parameters by first choosing a homomorphic computation, plaintext modulus, and a security parameter. This leads to a ring dimension and modulus $(n,Q)$ which 1) satisfies the security requirement and 2) allows for enough noise budget to allow decryption after homomorphic computation. Larger $Q$ gives more noise budget, but it also necessitates a larger dimension $n$ for security. The ratio of the ring dimension to the modulus bits, $n/\log_{2}Q$, stays roughly constant for a fixed security level.
Vectorized HE.
Vectorized plaintext spaces are used via polynomial interpolation: $X^{n}+1\mod t$ splits when $t$ is a prime satisfying $t=1\mod 2n$. In the vectorized case, each ciphertext encrypts a vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$ and the homomorphic operations are single-instruction multiple-data (SIMD) instructions. Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
SHE operations and costs.
BFV’s $\mathsf{Eval}$ calls five subroutines:
  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$.
  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$
  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$.
  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$.
  • $\mathsf{CtRotate}(\mathsf{ct},r)$ for $r\in[0,n/2)$ returns a ciphertext encrypting
    (((( $\mathsf{Rot}(\mathbf{v}),\mathsf{Rot}(\mathbf{\bar{v}}))\in\mathbb{Z}_{p}^{n/2}\times\mathbb{Z}_{p}^{n/2}$
    where $\mathsf{Rot}(\mathbf{v})$ ($\mathsf{Rot}(\mathbf{\bar{v}})$) is $\mathbf{v}$ ($\mathbf{\bar{v}}$) cyclically shifted to the left by r positions, if $\mathsf{ct}$ originally encrypted $\mathbf{v}^{\prime}=(\mathbf{v},\mathbf{\bar{v}})$. We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
Ciphertext rotation and ciphertext multiplication require an evaluation key since the secret changes with each rotation and ciphertext multiplication. This operation is called “key-switching” for rotations and “relinearization” for ciphertext multiplications.
Table 1 has concrete run times and noise growth for BFV. In general, addition and plaintext multiplication are efficient and addition has minimal noise growth. The latter has larger growth proportional to $t$. Ciphertext-ciphertext multiplication has large noise growth and is somewhat inefficient since it requires multiplying the ciphertext polynomials over the integers then rounding down back to $\mod Q$. Ciphertext rotation adds little noise but it is somewhat slow due to key-switching. Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
Operation Time (ms) Noise added (bits)
$\mathsf{CtCtAdd}$ $070$absent0$\approx$ 0≈ 0
$\mathsf{PtCtMult}$ $0.09$ $22$
$\mathsf{CtCtMult}$ $12.1$ $20$
$\mathsf{CtRotate}$ $3.6$ $\approx 0$
Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation. The polynomial degree $n$ is $8192$, the ciphertext modulus $Q$ has $150$ bits, and the plaintext modulus $t$ has $20$ bits. Time costs are measured with the SEAL library [CLP17] version 3.7.2 on a single core in a AWS t2.2xlarge instance.
Differential privacy
We use the standard notion of differential privacy (DP) [DMNS06, DKM+06]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$. Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
$\log\left(\frac{\Pr[M(X)=\zeta]}{\Pr[M(X^{\prime})=\zeta]}\right).$
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [Vad17, Lemma 2.3].
Lemma 1.
If $M_{1}(X),\dots,M_{k}(X)$ are each $(\epsilon,\delta)$-DP with independently sampled randomness, then $(M_{1}(X),\dots,M_{k}(X))$ is $(k\epsilon,k\delta)$-DP.
Anonymous networks.
Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [GRS99], or mixnets [Cha81]. Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
Private information retreival.
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [CGN98]. In this paper, we focus on single-server PIR protocols [KO97] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [ACLS18] and [ALP+21]. There is a notion of differentially-private PIR in the multi-server setting [TDG16, AIVG22], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
3 System Design
Figure 1: High-level overview of Wally. Throughout an epoch, clients sends encrypted real and fake queries at random time slots to an anonymization network which “washes” queries of identifying information.
A high-level design of Wally is in Figure 1. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Wally does not require any synchronization across clients. Each client can decide to participate or skip an epoch. However, we require a minimum of clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, privacy loss is the sum of the loss due to each epoch in which the client has participated. The correctness guarantee is that all clients’ queries are processed by the end of the epoch. Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a can participate at max $M$ epochs.
Each epoch is divided into second-long time slots, and privacy is guaranteed by the following:
  • Hiding the real queries of clients among fake queries (differential privacy).
  • Hiding the content of real and fake queries (somewhat homomorphic encryption).
  • Removing the identifiable information from each query and shuffling the batch of queries received within a second (anonymization network).
  • Hiding the information revealed by queries order and time of arrival (random query schedule)
Within each epoch, Wally performs a semantic search that outputs the most relevant results to client queries, similar to insecure search engines.
Embedding-based search.
Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, it is used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [DCLT19].
In the following sections, we introduce the insecure version of Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
3.1 Insecure Search Pipeline
Once the client query and the server index are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
  • Server initialization. At initialization, the server divides the entries into $K$ clusters using $K$-means clustering. A heuristic-based technique partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$ and outputs these clusters with their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
  • Pick nearest centroid. To make a query, the client first finds the cluster nearest to the query embedding. For this, the client computes a cosine similarity between the query and each centroid embedding and picks a cluster whose centroid has high cosine similarity.
  • Server computation. The client then sends the nearest cluster ID $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes cosine similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
  • Pick nearest entry. The client sorts these scores and picks the entry closest to the query.
  • Metadata retrieval. The client then fetches the metadata associated with the closest entry from the server.
The clustering technique provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is in a different cluster than the cluster with the closest centroid. Therefore, finding a balance is important for the overall usability of the system.
The above protocol is not private. To ensure that the protocol does not reveal any information about the client, it must hide the following information from the server:
[…]
  • Nearest cluster. The client reveals to the server the nearest cluster id. Recall that clustering puts highly correlated documents together in a cluster. Therefore, cluster id reveals the kind of data the client holds. For example, if the client query is an image, the cluster ID might reveal whether the image is a landmark, e.g., a castle or a bridge.
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about client queries.
We show how we hide the above data below and in Section 5.
Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
3.2 Hiding the Query Embedding with SHE
To hide a query embedding, we rely on vectorized, RLWE-based SHE introduced in Section 2. The client query is an encrypted vector, and the server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
Server database.
The server divides the database, of size $N$, into $K$ clusters each with at most $N^{\prime}$ embeddings. For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size. Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d*d$, and each column within a slice is an embedding. Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$. For now, we assume that each diagonal of a slice is a separate plaintext vector of size $n=d$. That is each slice is encoded as $d$ plaintexts vectors and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$. This means each plaintext vector consists of a single index from each embedding in a slice in increments of one. The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
Dot product computation.
Recall, from Section 2, for cosine similarity, the server needs to compute a dot-product between two embeddings. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $d=4$. This diagonal method is to minimize homomorphic rotations [HS14, HS18].
Database and query packing.
In the above description, we assume that the polynomial dimension $n$ equals embedding dimension $d$. However, $n$ is often significantly larger than $d$. In this setting, each plaintext vector could hold multiple diagonals. Our scheme packs diagonals across the slices in a single plaintext vector. Concretely, the $j$-th plaintext vector in a cube holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices. In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$. Therefore, by using packing, the entire cube can be encoded with only $d$ plaintext vectors.
The client query encrypts a plaintext vector of $\lfloor{n/d}\rfloor$ repetitions of the query embedding. This query still allows the server to align the entries of query embeddings by homomorphically rotating left in increments of one. The number of ciphertext-plaintext multiplications is now $\lceil N^{\prime}/n\rceil*d$ while the response consists of $\lceil N^{\prime}/n\rceil$ ciphertexts. Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
Recall from Section 2 that homomorphic rotation is a costly operation. Even after packing, the server still performs $d$ rotations. Looking ahead to Section 6, we explain how we reduce the number of rotations to $2\sqrt{d}$ by using the standard the baby-step giant-step optimization [HS18].
Float-to-integer encoding.
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. Wally converts each float into an integer by multiplying with the appropriate scaling factor and only taking the integer part of the result. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding nearest cluster is a significant performance bottleneck in state-of-the-art schemes like Tiptoe. Specifically, to hide the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database. Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the strongest possible sense, but this results in a prohibitively high server computation.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is that the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic, obliviousness Tiptoe achieves.
The starting point of our solution is that in large-scale systems, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism allowing clients to hide their queries among the crowd of all queries. Our proposed approach is based on $(\epsilon,\delta)$-differential privacy. The approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
Intuitively, one would expect that routing all the traffic through an anonymization network would provide sufficient privacy as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can potentially learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known to not access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
Our protocol randomizes queries to guarantee differential privacy. Specifically, we make the following two changes:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks on of the clusters.
  • Random query schedule. The client makes queries at random instances within the epoch. For this, we assume that each epoch is divided into a slots, each one second long. For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
In Section 5, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client might want to retrieve metadata associated with it. When each entry has a small amount of metadata, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally picks the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run PIR to fetch the metadata of only the most relevant entry.
3.5 Keyword Search
Wally also supports keyword search in which the client is interested in retrieving data associated with a keyword. Using a hash function, the server database is divided into random $K^{\prime}$ buckets (instead of clusters). For each query, the client locally finds the bucket in which the keyword falls and then uses the keyword PIR scheme to fetch a specific record within a bucket.
Wally uses a keyword PIR construction based on cuckoo hashing and index PIR. At a high level, at initialization, the server uses the cuckoo hashing to map the $N$ database elements into a table of size $O(N)$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [ALP+21] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
Each bucket consisting of $N/K^{\prime}$ entries represented as $\sqrt{N/K^{\prime}}\times\sqrt{N/K^{\prime}}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.
  1. 1.
    First expand $\hat{q}$ to two $\sqrt{N/K^{\prime}}$-length ciphertext vectors. Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$. To expand, the server uses the oblivious expansion algorithm given in [ALP+21]. This algorithm requires $2\sqrt{N/K^{\prime}}$ homomorphic substitutions.
  2. 2.
    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/K^{\prime}$ plaintext-ciphertext multiplications and $N/K^{\prime}-\sqrt{N/K^{\prime}}$ ciphertext additions.
  3. 3.
    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N/K^{\prime}}$ ciphertext-ciphertext multiplications and $\sqrt{N/K^{\prime}}-1$ ciphertext additions. to compute.
The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule. Section 6 discusses various optimizations we implemented for MulPIR. Overall, our optimizations result in around a three times reduction in server computation and a two-three times reduction in request size.
[…]
Figure 4: Simplified protocol for Wally. Each client generates queries with Algorithm 2 and sends them to the anonymization network, $\mathcal{AN}$, according to the randomized schedule described in Algorithm 4. Cluster $j$ receives anonymous queries denoted as $(\tilde{q}^{j}_{i})$ and computes each response $(r^{j}_{i})$ via SHE.
Algorithm 1 $\mathsf{ServerInit}$
Input: Semantic search database $\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]}$, Number of clusters $K$
Output: Processed semantic $D_{\text{sem}}$,
Clusters centroids $c$.
1:$(\mathbf{c},\mathbf{C})\leftarrow\textsf{K-MeansClustering}(\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]},K)$triangleright dividing semantic data into $K$ clusters
2:for $i\in[K]$ do
3:     for $j\in[d]$ do $\triangleright$ Diagonal packing
4:         for $l\in[|C[i]|/d]$ do
5:              $\mathbf{p}[{j}]=\mathbf{p}[{j}]||[C[i,0]_{l*C[i]/d+j}^{j},$
6:
7:              $\cdots,C[i,0]_{l*C[i]/d+j+d-1}^{d-1}]$
8:         end for
9:         $\mathbf{p}[j]=\mathsf{EncodePtVec}(\mathbf{p}[j])$
10:     end for
11:     for $l\in[h]$ do $\triangleright$ Rotations for BSGS
12:         for $j\in[g]$ do
13:              $\mathbf{p}[gl+j]=\textsf{PtxtRotate}(\mathbf{p}[gl+j],-gk)$
14:         end for
15:     end for
16:     $D_{\text{sem}}[i]=p$
17:end for
18:Output $D_{\text{sem}},c$
4 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch with $U$ clients.
Server initialization.
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm first divides the input embeddings into $K$ clusters using K-means clustering. We pick $K$ such that the maximum cluster size is less than SHE polynomial dimension $n$ with high probability, which ensures that for each cluster all the embeddings can be encoded into $d$ plaintext vectors. The algorithm then iterates over each cluster separately. Recall each cluster consists of $|C[i]|/d$ column-wise $d*d$ matrices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts and rotating each group in increasing multiples of $-g$. Note that these rotations are performed on plaintexts so their cost are negligible.
Query generation.
The client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally stores the set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$.
The algorithm, then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries. To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$. In other words, the schedule $\mathbf{S}$ maps slot indices to the query indices. At a particular time slot $t$ within an epoch, the client will make queries (independently) indicated by $S[t]$.
Server computation.
The server uses Algorithm 4 to generate a response for every query received. The algorithm first aligns the query ciphertext $\mathsf{ct}$. That is, first it copies the query ciphertext $g$ times and then rotates each copy left in increments of one, in total $g$ rotations. Then, the algorithm performs a dot-product between rotated query ciphertexts and the plaintext vectors for cluster $\mathsf{id}$. Recall that we ensure each cluster is encoded into $d$ plaintexts. The algorithm iterates in groups of $g$ plaintexts; within each group, multiply $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts. These iterations yield $h=\sqrt{d}$ ciphertexts. After that, the algorithm iterates over the $h$ resulting ciphertexts, rotating right each in increments of $g$ and summing all of them into a single ciphertext $R$. This step requires a total of $h$ rotations. The algorithm returns $R$ and cluster $\mathsf{id}$’s metadata.
Algorithm 2 $\mathsf{ClientQuery}$
Input: client queries $q=\{q_{i}\}_{i\in\Delta}$
Clusters centroids $c=\{\mathbf{c}_{l}\}_{l\in[K]}$
$U$ is the number of clients in the system
Output: Real and fake queries list $\mathbf{Q}$.
1:for $l\in[\Delta]$ do $\triangleright$ generate real queries
2:     $\mathbf{p}[1:N/K]=\mathbf{e}_{l}||\mathbf{e}_{l}||\cdots||\mathbf{e}_{l}$
3:     $(\mathsf{sk}_{l},\mathsf{evk}_{l})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
4:     $\mathsf{ct}_{l}\leftarrow\mathsf{EncSemPt}({\mathsf{sk}_{l}},\mathbf{p})$
5:     Pick cluster $\mathsf{id}_{l}$ nearest to $\mathbf{e}_{l}$ from $c$
6:     $Q\leftarrow Q||(\mathsf{ct}_{l},\mathsf{evk}_{l},\mathsf{id}_{l})$
7:end for
8:for $i\in[K]$ do
9:     $F_{i}\leftarrow\mathcal{NB}(r/U,p)$triangleright number of fake queries
10:     for $j\in[F_{i}]$ do $\triangleright$ generate fake queries
11:         $(\mathsf{sk}^{\prime}_{j},\mathsf{evk}^{\prime}_{j})\leftarrow\mathsf{KeyGen}(1^{\lambda})$
12:         $\mathsf{ct}_{j}^{\prime}\leftarrow\mathsf{EncSemPt}(\mathsf{sk}_{j}^{\prime},\mathbf{0})$
13:         Pick random cluster $\mathsf{id}_{j}$ from $c$
14:         $Q\leftarrow Q||(\mathsf{ct}_{j}^{\prime},\mathsf{evk}_{j}^{\prime},\mathsf{id}_{j})$
15:     end for
16:end for
17:Randomly permute $Q$
18:$\mathbf{S}\leftarrow\mathsf{RandScheduleGen}(Q)$triangleright generate schedule
Algorithm 3 $\mathsf{RandScheduleGen}$
Input: List $\mathbf{Q}$ of real and fake queries
1:Epoch length is $T$
2:Initialize a schedule $\mathbf{S}$triangleright $T$ empty lists
3:for $j\in[|\mathbf{Q}|]$ do
4:     $i\leftarrow[T]$triangleright Sample random slot
5:     $\mathbf{S}[i]\leftarrow\mathbf{S}[i]||j$
6:end for
7:return $\mathbf{S}$
Algorithm 4 $\mathsf{ServerComputation}$
Input: Query $(\mathsf{ct},\mathsf{evk},\mathsf{id})$
Output: Encrypted scores and data $r$.
1:for $j\in[g]$ do $\triangleright$ Align query
2:     $\mathsf{ct}^{\prime}[j]=\mathsf{SHERotate}_{\mathsf{evk}}(\mathsf{ct},j)$
3:end for
4:$D=D_{\text{sem}}[\mathsf{id}]$triangleright Dot-Product
5:$R=\sum_{k\in[h]}\mathsf{SHERotate}_{\mathsf{evk}}(\sum_{j\in[g]}D[j+kg]*\mathsf{ct}^{\prime}[j],gk)$
6:$\textsf{data}\leftarrow\textsf{data}_{\mathsf{id}}$triangleright Cluster $\mathsf{id}$’s metadata
7:return $r:=(R,\textsf{data})$
5 Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
We prove this argument in two steps: First, we show that a noisy histogram output by a curator, which gets users’ queries as input and uses negative binomial $\mathcal{NB}$ mechanism to sample fake queries, is $(2\epsilon,2\Delta\delta)$-DP in the central model. Second, we show that the view of the server in an epoch can be simulated only using the curator’s output.
5.1 DP Security in the Central Model
We define the central model as follows:
Central model.
The curator performs the following:
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$, where $U$ is the total number of clients, and $R_{u}$ is each client’s set of queries and will be at most $\Delta$. Let $R$ denote the list of all real queries.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathcal{NB}(r,p)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real and fake queries, randomly permute them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$.
  4. 4.
    Send $\mathsf{list}$ to the server.
Another way to view it is that the curator sends a noisy histogram of queries to each cluster to the server. Recall that each client can contribute at most $\Delta$ queries.
We then prove the following theorem:
Theorem 1.
For any query database of size $N$, number of clusters $K$, differential privacy parameters $\epsilon,\delta\in(0,1)$, and $\Delta\in\mathbb{N}$, let $p=e^{-0.2\epsilon/\Delta}$ and $r=3(1+\log(1/\delta))$. Then, the negative binomial mechanism $\mathcal{NB}(r,p)$ is $(2\epsilon,2\Delta\delta)$-DP for $\Delta$-histogram.
Proof.
We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram. We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing a client with another.
For a particular cluster $b$ in the histogram, we define $M_{b}(\cdot):=R_{b}+\mathcal{NB}(r,p)$, where $R_{b}$ is the total real queries for the cluster $b$ in the database. Note that this is the curator’s exact output for each cluster. We break down the proof into two edge cases: 1) the simpler case where $X$ and $X^{\prime}$ differ by a client which sends all their messages to a single cluster and 2) they differ by a client which sends $\Delta$ messages to $\Delta$ different clusters.
Simple case. Consider two neighboring databases $X$ and $X^{\prime}$ which differ in queries contributed by a single client. Concretely, in $X$ the client contributes $\Delta$ queries to a cluster $b$ and in $X^{\prime}$ instead contributes to another cluster $b^{\prime}\not=b$. Note that when $j\notin\{b,b^{\prime}\}$ then $M_{j}(R)$ and $M_{j}(R^{\prime})$ are identically distributed. Thus, the privacy provided by $M_{j}$ is 0-DP for all $j\notin\{b,b^{\prime}\}$.
Now consider the cluster $b$; the two databases differ by $\Delta$ real queries for this cluster. By the following, lemma $M_{b}$ provides $(\epsilon,\delta)$ privacy on cluster $b$.
Lemma 2 (Theorem 13 [GKMP20]).
[…]
We can make a similar argument for a cluster $b^{\prime}$. Therefore, by basic composition, Lemma 1, the total privacy provided by $M$ for these neighboring databases is $(2\epsilon,2\delta)$—DP.
General case. Now, we consider a general case where two neighboring databases that differ in queries from a single client. In database $X$ the client contributes queries $\gamma_{j}$ for cluster $j$ and in database $X^{\prime}$ the client contributes queries $\gamma_{j}^{\prime}$ for the cluster $j$. Define $\Delta_{j}:=|\gamma_{j}-\gamma_{j}^{\prime}|$. The total difference between two neighboring databases is given by $\sum_{j}^{B}\Delta_{j}\leq 2\Delta$.
Let $\epsilon_{j}:=\frac{\epsilon\Delta_{j}}{\Delta}\leq\epsilon$. Further, we only consider the clusters where $\Delta_{j}>0$ since the clusters where $\Delta_{j}=0$ are already 00-DP. (Equivalently, $\epsilon_{j}=0$ for these $j$.) Then, Lemma 2 implies cluster $j$ is $(\epsilon_{j},\delta)$-DP for $\Delta_{j}$-summation for the same mechanism since
$\frac{\epsilon}{\Delta}=\frac{\epsilon}{\Delta}\cdot\frac{\Delta_{j}}{\Delta_{j}}=\frac{\epsilon_{j}}{\Delta_{j}}.$
Simple composition over all the clusters yields
$\sum_{j}\epsilon_{j}=\frac{\epsilon}{\Delta}\sum_{j}\Delta_{j}\leq 2\epsilon.$
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ.
We emphasize that the proof is over a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
5.2 Generating view of the server
We show how the curator’s output is sufficient to simulate the server’s view generated by distributed DP algorithm defined in Algorithm 2 together with an anonymization network (AN). We also require pseudorandom ciphertexts and evaluation keys, a common assumption for RLWE-based SHE schemes. The latter is also why we cannot reuse evaluation keys.
Theorem 2.
Let $\mathcal{AN}$ be an anonymous network that 1) removes all identifying information from messages it receives and 2) randomly permutes all the messages it receives over a second-long slot. Let each honest client run the distributed mechanism in Algorithm 2 with messages sent to the server via $\mathcal{AN}$. Then, the view of the server in an epoch can be simulated using the curator noisy histogram output described above in the central model.
Proof.
We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard hybrid argument in cryptography.
Hybrid 0: Every client is the same as the real client given in Algorithm 2 except it does not include encrypted embeddings with each request (real or fake). Hybrid 0 is equivalent to the real client finding the nearest centroid for each real query and a fake random centroid for each fake query. The client then picks a random slot for each query and send associated queries at each slot.
Hybrid 1: In this hybrid, we replace all the clients with a single simulator. The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ is real queries for $u$-th client. The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$. From the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
Hybrid 2: In this Hybrid, the simulator’s input is the same as in the previous Hybrid 1. The simulator is defined as follows.
  1. 1.
    Collects the real queries from all the clients, $\{R_{u}\}_{u\in[U]}$. Call it a list of real queries $R$.
  2. 2.
    For each cluster $b$, samples the number of fake queries $F_{j}\leftarrow\mathsf{NB}(p,r)$. Add it to the fake queries list $F$.
  3. 3.
    Given the list of real queries $R$ and fake queries $F$, randomly permutes them: $\mathsf{list}\leftarrow\mathsf{permute}(R,F)$
  4. 4.
    Initialize schedule list $S$ of size $z$ (where z is the total number of slots in an epoch). For each query $q\in\mathsf{list}$, pick a random slot $p\leftarrow z$ and set $S[p]=S[p]\cup q$.
  5. 5.
    For each slot $p\in S$ send queries in $S[p]$ to the server.
There are two differences between Hybrid 2 and Hybrid 1:
  1. 1.
    In Hybrid 1, each simulated client samples fake queries for each cluster from a distribution $\mathcal{NB}(r/U,p)$. While in Hybrid 2 (Line 2) the simulator samples fake queries per cluster from $\mathcal{NB}(r,p)$.
  2. 2.
    In Hybrid 1, each simulated client generates a schedule for its queries independently. While in Hybrid 2, the simulator generates a schedule for queries.
To show that Hybrid 2 and 1 are indistinguishable, we use the definition of infinitely divisible distribution, Definition 1.
Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)\ \sum_{i=1}^{U}\mathcal{NB}(r/U,p)$. Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1. Observe that a random slot is picked for each query independently in both Hybrids. Therefore, the query schedule generated in both Hybrids is indistinguishable.
To complete the proof, note that in Line 3, the simulator generates $\mathsf{list}$ in the same way as the curator in the central model. After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
6 Optimizations
Here we present our optimizations for semantic and keyword search. Our optimizations revolve around the [Bra12, FV12] and BGV [BGV14] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [ALP+21]. First, we describe the standard RNS optimization in SHE [GHS12] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [HLHD22].
RNS form.
Often $Q$ is larger than machine-size words and we use $Q$ as the product of machine-sized primes $Q=q_{0}q_{1}\cdots q_{l}$. This allows us to store a ciphertext as two polynomials each stored as an $n\times(l+1)$ matrix of integers by the Chinese remainder theorem (CRT) on $Q$. Further, we require each $q_{i}$ to be NTT friendly, $q_{i}=1\mod 2n$. This allows us to use the number theoretic transform (NTT) to efficiently switch between each evaluation form and coefficient form in time $O(n\log_{2}n)$, since the NTT is a modular version of the fast Fourier transform over a prime modulus. We can efficiently compress a ciphertext by modulus-switching down to a smaller modulus, $Q^{\prime}$ where $Q^{\prime}|Q$, e.g., $Q^{\prime}=q_{0}$, without affecting the ciphertext’s noise budget.
Dropping ciphertext LSBs.
We use Cheetah’s [HLHD22] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [HHC+23]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
$2^{b_{0}}c_{0}^{h}+2^{b_{1}}c_{1}^{h}s=\lfloor Q/t\rceil m+e+e_{drop}$
for
$e_{drop}=-c_{0}^{l}-c_{1}^{l}s.$
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined our BFV plaintext CRT optimization described in the next subsection.
6.1 Approximate Nearest Neighbors Optimizations
The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
Minimal rotation keys and bandwidth.
Wally uses the “baby-step giant-step” (BSGS) optimization [HS18] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [HS18] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [GHS12] key-switching strategy described in-detail by Kim et al. [KPZ21].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [APS15] [^4] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
Re-using BFV evaluation keys via the plaintext CRT.
Wally achieves minimal bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows us to keep a set amount of plaintext bits while re-using one set of evaluation keys across all plaintext spaces since BFV’s evaluation keys are independent of the ciphertext modulus, unlike BGV evaluation keys. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{i}$ for NTT-friendly primes $\{t_{i}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}\cdots t_{l}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$. Therefore, we can drop the ring dimension $n^{\prime}=4096$ and use two ciphertexts, modulo $t_{0}=40961\approx 2^{15}$ and modulo $t_{1}=65537\approx 2^{16}$ for $t_{0}t_{1}\approx 2^{31}$, half the bandwidth from ciphertexts, and use evaluation keys in dimension $n^{\prime}$, or one fourth the bandwidth from evaluation keys. Concretely, this leaves us with $113$kB for ciphertexts and $170$kB for evaluation keys, $283$kB in total, over a 3×3$\times$3 × improvement in query size without sacrificing plaintext precision.
In general, the plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: six polynomials versus 2 polynomials. This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching. Another advantage of the plaintext CRT is that it allows for very high precision computations which do not wrap around mod $t$. One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
6.2 PIR Optimizations
Uneven dimensions.
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, SEAL [CLP17] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [MR23].
Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$. For any fixed HE parameter set, the server’s compute time is dominated by ciphertext multiplications and rotations and that the former is a constant factor $\gamma$ more than the latter. This simplification yields a compute time of $2d_{1}t_{r}+2d_{2}t_{r}+d_{2}t_{\times}$ where $t_{\times}$ and $t_{r}$ are the times to multiply and rotate ciphertexts, respectively. More simply, the total compute time, measured in terms of rotations, is
$(\gamma+2)d_{1}+2d_{2}=(\gamma+2)C/d_{2}+2d_{2}$
rotations. Minimizing this as a function in $d_{2}$ gives $d_{2}=\sqrt{1+\gamma/2}\sqrt{C}$.
This optimization saw an improvement of $30-40\%$ in MulPIR run times. We chose this concrete analysis since SHE parameter regimes are restricted per use case and it is much simpler than an asymptotic analysis.
Linearizing parts of query expansion
MulPIR is an optimized query expansion which originated with Angel et al. [ACLS18]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
Lazy rescaling in BFV multiplication.
We apply the lazy rescaling technique of Kim et al [KPZ21] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [KPZ21, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Handling Very Large Entries
In MulPIR the database is represented as a plaintext matrix $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$. Conventionally, the response is computed as
[…]
This approach results in $n\cdot t$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large. But it will blow up with extremely large entries (e.g. when they are high-resolution photos).
To overcome this challenge, we can swap the order of the computation as
[…]
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size and $t$. This optimization reduce the computation overhead when $t$ is larger than the first dimension $d_{1}$.
Keyword PIR.
MulPIR’s keyword PIR functionality is given by using Cuckoo hashing to find the entry’s proper index. We noticed for our cluster sizes $C$, we are able to split the cluster into two hash tables and perform $1$-hash Cuckoo hashing on each without increasing the (empirical) failure probability. This resulted in a 2×2$\times$2 × improvement in query size and server computation time.
Low-level SHE Optimizations
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [BEHZ16] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
We add a conditional, lazy modular reduce to Harvey’s NTT [Har14]. The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.
7 Performance
In this section, we demonstrate Wally’s scalability and concrete performance as seen in experiments. We also track Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
Experimental setup.
We ran server-side computations on an AWS c5.24xlarge instance with 192GB of RAM, single-threaded. We extrapolate throughput from the single-threaded times seen on this AWS instance.
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Exp. query (MB) 0.67 0.84 16.2
Exp. response (MB) 0.16 0.18 3.2
Queries/s 2998 1499 47
8M queries (hrs) 0.4 1.5 47.5
Fake query 3×3$\times$3 × 3×3$\times$3 × --
overhead
Precision (bits) 6 13 4
Privacy DP DP SHE
Network req. AN AN --
Table 2: Wally performance for a parameter set with BFV dimension $n=4096$ on a 35M entry dataset with the maximum cluster size as $4096$ compared to Tiptoe. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8M real queries, or 24M in total. All Wally responses are computed as the encrypted scores plus $44$kB of metadata.
Table 2 describes our main experiments in comparison with Tiptoe. Here, we use two parameter sets for Wally: first, we use a small BFV parameter set ($n=4096$, $\log_{2}t=15$, and $\log_{2}q=83$) and second, a plaintext CRT (Section 6) parameter set which includes the former. The plaintext precision is $(\log_{2}t)/2-(\log_{2}d)/4$ since the inner-product computation grows heuristically as $\sqrt{d}\|\mathbf{\tilde{e}}\|_{\infty}^{2}$ where $\mathbf{\tilde{e}}=\alpha\cdot\mathbf{e}$ is a scaled (unit) embedding vector, $\|\mathbf{e}\|_{2}=1$. The former parameter set enables six bits of plaintext precision whereas the latter enables 13 bits.
We set $\Delta=2$ messages per client per epoch and calculated the DP overhead by using $(\epsilon,\delta)=(1,2^{-20})$ after the loss in Theorem 2 with four million clients per epoch. The fake query overhead is derived by the expected number of fake queries, for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
where $K$ is the total number of clusters and $U$ is the total number of clients.
Wally’s ability to cluster the dataset and keep server computation time proportional to the maximum size of a cluster is demonstrated in its small per-query server computation time, $95$ms for each real or fake query, and its relatively small expected query size, $679$kB. We keep each cluster size less than the total number of plaintext slots in each BFV ciphertext, $n=4096$. Otherwise, Wally can compute similarity scores in parallel if the cluster size is larger than $n$. Large meta-data sizes can be retrieved by adding another round of PIR after the similarity computation. The throughput times represent how many queries per second can be processed on $K=8545$ clusters during an epoch needed to serve $24$M total queries, real and fake. This throughput calculation leads to a minimum epoch size of $44$ minutes serving roughly 3k queries per second. This compares to the $47.5$ hours Tiptoe would take to serve $8$M queries.
Comparing to Tiptoe.
Wally has the following two extra assumptions compared to Tiptoe:
  • Wally requires sufficiently many honest clients in the system during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short amount of time, e.g., one second time slots.
In other words, Wally trades off scalability for privacy. Such a tradeoff is needed since oblivious, cryptographically secure systems require the server to compute over the entire database and cannot scale to many clients.
Tiptoe ran its encrypted similarity computation on a cluster of $80$ r5.xlarge AWS instances, each with 4 vCPUs and 32 GB of DRAM, and saw a latency of $11.2$ seconds per query, of which is a significant amount of per-query preprocessing (8.7 seconds). Each server held $8$GB to $12$GB of the plaintext database. We emphasize that each Tiptoe query computes of the entire database (all clusters), and we estimate their total single threaded time as 80×80$\times$80 × their per query time, or $183$ seconds per query. We obtained this estimate by dividing the pre-processing (token) times and the download bandwidths in Table 7 in [HDCZ23] by six since our embedding dimension is 3×3$\times$3 × higher and Tiptoe generates two pre-processing elements per query, “tokens”: one each for a similarity computation and PIR. The non-pre-processing times and bandwidths, as well as the token upload bandwidth, are divided by 3×3$\times$3 × again due to the difference in embedding dimension and that they use the same token upload for both the similarity computation and PIR. Next, the server compute times are multiplied by $80$ since they use $80$ clusters, yielding an estimated $163$ seconds, single-threaded, for a homomorphic similarity computation, i.e., a Wally query. Given this time, we divide it by the number of clusters we expect in Wally, $8545$, to estimate throughput in Table 2.
In comparison, Wally can process each query in parallel since queries, real and fake, have their computation restricted to a single cluster. If there are $U$ clients making at most $\Delta$ queries, then Wally can set a epoch time as the minimal amount of time needed to compute $\phi\times U\times\Delta$ queries. Since each client sends a query at a random time over an epoch, the average response time is half the epoch whereas the worst case time is the epoch length. We derived Wally’s throughput numbers in Table 2 using $95$ms per query, $U=4$M, $\Delta=2$, and $\phi=2$. This leads to $24$M queries in 44 minutes, of which $8$M are real queries, on $K=8545$ clusters.
Scaling DP with $\epsilon$.
Figure 5 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients. This is derived by dividing Equation 1 by the number of clusters $K$. Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
Figure 5: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section 6. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 95 190 580
Table 3: Here we list the query and response size per each query (real and fake). Each response size includes $44$kB of metadata. Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
Scaling with the number of clients.
Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system. We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients. As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads. Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
Number of Clients 0.5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$65.65.65.65 $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$61.61.61.61
Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
7.1 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an M1 Max laptop. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters. The request size and the response size for each entry in the table is 728KB and 39KB, respectively.
Database Entries Entry Size (bytes) Time (s)
100K 10 0.2
100K 100 0.5
1M 10 0.7
1M 100 2
10M 10 2.7
100M 100 6.3
Table 5: Keyword-PIR, server run times for Wally. These times include the optimizations in Section 6.
8 Conclusion
Wally demonstrates that private search systems can be scalable to systems with many users. Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters. Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases. This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.

### v3

@@
-Scalable Private Search with Wally
+Wally: An Efficient Private Search Engine
 Abstract
-This paper presents Wally  a private search system that supports efficient semantic and keyword search queries against large databases.
-When sufficient clients are making the queries, Wally performance is significantly better than previous systems.
+This paper presents Wally, a private search system that supports efficient semantic and keyword search queries against large databases.
+When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems.
 In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry.
 As a result, performance degraded proportionally with the number of entries in the database.
-In Wally we get rid of this limitation.
+In Wally, we get rid of this limitation.
 Specifically, for each query the server performs cryptographic operations only against a few database entries.
-We achieve these results by requiring each client to add a few fake queries, and sends each query via an anonymous network to the server at independently chosen random instants.
-Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake, Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
+We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants.
+Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake.
+Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
 The number of fake queries each client makes depends inversely on the number of clients making queries.
 Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases.
-Concretely, Wally can serve $8$M requests at a rate of 3,000 queries per second.
-That is around 60x higher than the state-of-the-art scheme.
+Concretely, Wally can process eight million queries in just 39 mins.
+That is around four orders of magnitude less than the state of the art.
 1 Introduction
@@
 These search engines improve the quality of our lives by promptly providing helpful information.
-Nevertheless, they require learning the client’s query to respond accurately to each query, hence providing no query privacy.
+Nevertheless, they require learning the client’s query in order to respond accurately, hence providing no query privacy.
 In some situations, learning the query could reveal sensitive and personal information about the client.
@@
 Private search engines opt to provide query privacy.
-However, these systems have significantly higher computation and communication overhead than the state-of-the-art insecure solutions.
-For example, in Tiptoe [HDCZ23], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries.
-The high computation is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in.
+However, these systems have significantly higher computational and communication overhead than the state-of-the-art insecure solutions.
+For example, in Tiptoe [29], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries.
+The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in.
 Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick.
 We detail this trade off in a later section..
-Additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
-In this paper, we present Wally, a privacy preserving search engine that efficiently scales to performance constraints of large-scale search engines.
+An additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
+In this paper, we present Wally, a privacy preserving search engine that efficiently scales to the performance constraints of large-scale search engines.
 Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword.
@@
 We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries.
-In our protocol, the server computation and the communication get amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
+In our protocol, the server computation and the communication are amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
 Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy.
-In Wally the server learns noised differentially private histogram of anonymized queries from all clients.
-We emphasize that Tiptoe, is fully oblivious, i.e., the server learns nothing besides that a particular client is making a query.
-Therefore, the privacy guarantee of Tiptoe is stronger than Wally  however differential privacy is an accepted standard for strong privacy [DR+14].
-Specifically, in Wally every client is guaranteed that all the statistical inferences are insensitive to their data.
+In Wally, the server learns a noised differentially private histogram of anonymized queries from all clients, over a partition of the server database.
+We emphasize that previous works, e.g., Tiptoe, are fully oblivious.
+That is, the server learns nothing besides that a particular client is making a query.
+Therefore, the privacy guarantee of schemes like Tiptoe is stronger than Wally.
+However, differential privacy is an accepted standard for strong privacy [19].
+Specifically, every client in Wally is guaranteed that all the statistical inferences are insensitive to their data.
 In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
@@
 Within each epoch, a fixed number of users make queries.
-The epoch size is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch.
-The server database is divided into $C$ clusters.
+The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch.
+Server side, the database is divided into $K$ clusters.
 Each query consists of a relevant cluster index, and the server response is all the entries in the cluster.
@@
 To prevent this leakage, Wally implements the following two changes at the client.
-To ensure that the server does not learn information by observing the access patterns, in Wally each client also makes fake queries.
+First, to ensure that the server does not learn information by observing the access patterns, each client in Wally makes fake queries.
 For each fake query, the client picks a randomly selected bucket.
@@
 We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
-Wally also ensures that arrival time and order of the queries leak no information about the client to the server.
+Second, Wally also ensures that arrival time and order of the queries leak no information about the client to the server.
 To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch.
@@
 For large clusters, this would result in very high response overhead.
-In Wally  we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead.
+In Wally, we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead.
 Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE.
@@
 We propose various optimizations that significantly improve performance for both kinds of searches.
+Overall, a database with 35 million entries Wally can process eight million real queries in just 39 minutes on 320 cores set up, a drastic improvement over the state-of-the-art, which would have required years to process the same queries.
+Additionally, the size of each request is 53 times smaller and the size of each responses is 3.5 times smaller compared to Tiptoe.
 Our contributions are as follows:
@@
   • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
+Due to space limitation, we discussed related works in Appendix A.
 Limitation.
-Wally requires decent number of clients always present in an epoch to make queries.
-We argue that this assumption aligns with current search engines, for example a hundred of millions of people make queries in Google search daily [Fle24].
-We remark that for systems with few users, overhead due to fake queries could get quite high, therefore for these systems fully oblivious protocols like Tiptoe will be suitable.
+Wally requires a set number of clients always present in an epoch to make queries.
+This assumption aligns with current search engines.
+For example, hundreds of millions of people make queries in Google search daily [22].
+We remark that for systems with few users, the overhead due to fake queries could get quite high, therefore fully oblivious protocols, like Tiptoe, for these systems would be more suitable.
 2 Background
@@
 That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
-2.1 Homomorphic encryption
-We use the BFV [Bra12, FV12] SHE scheme [^2] BFV is also a fully homomorphic scheme, which can homomorphically compute any function, with large enough parameter sets.
-We only focus on its simpler SHE parameter regime in our paper.
-where each plaintext is a polynomial in $\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and $t$ is a prime or prime power.
-Wally can also be implemented with the BGV [BGV14] SHE scheme with similar performance.
-Both schemes are post-quantum SHE schemes based on the ring-learning-with-errors (RLWE) problem [Reg05, LPR10].
-A BFV ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})\in R_{Q}^{2}$ reduced modulo $Q$ and $X^{n}+1$ for a power of two $n$, $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$.
-The number $n$ is a ring dimension, $n\in\{2^{12},2^{13}\}$ in our use cases, and $Q$ is a ciphertext modulus with $\log_{2}Q=83$ or $\log_{2}Q=155$ since we use 128-bit secure post-quantum security parameters [ADPS16].
-A BFV ciphertext satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=\lfloor Q/t\rceil m+e$ where $s$ is a secret key polynomial with small coefficients, in $\{0,\pm 1\}$, $e$ is small error polynomial, $t$ is the plaintext modulus, and $m\in\mathbb{Z}_{t}[X]/(X^{n}+1)$ is the plaintext polynomial.
-This error polynomial’s entries grow as we compute more homomorphic operations on the ciphertext and we can only decrypt if $\|e\|_{\infty}<Q/2t$.
+2.1 (Somewhat) Homomorphic Encryption
+Fully homomorphic encryption (FHE) is a special scheme allowing arbitrary computation over ciphertexts.
+FHE for arbitrary computation is still very expensive.
+To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports computation up to a fixed depth.
+We focus on SHE schemes based on the Ring Learning with Errors (RLWE) problem.
+Concretely, Wally can be implemented using any SHE scheme such as BFV [21] or BGV [11] that share the following structure.
+A plaintext $m$ is a polynomial in a ring $R_{t}=\mathbb{Z}_{t}[X]/(X^{n}+1)$ (with degree at most $n-1$) and plaintext modulus $t$ is a prime or prime power.
+The secret key $s$ is a polynomial of degree $n-1$ with small coefficients, in $\{0,\pm 1\}$.
+A ciphertext is a pair of polynomials $\mathsf{ct}=(c_{0},c_{1})=(a,as+m+e)\in R_{Q}^{2}$ where $Q$ is ciphertext modulus and $R_{Q}:=\mathbb{Z}_{Q}[X]/(X^{n}+1)$.
+Here $a$ is picked uniform randomly and $e$ is a noise polynomial with coefficients sampled from a bounded Gaussian distribution.
+A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$ This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$.
 We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
@@
 The ratio of the ring dimension to the modulus bits, $n/\log_{2}Q$, stays roughly constant for a fixed security level.
-Vectorized HE.
-Vectorized plaintext spaces are used via polynomial interpolation: $X^{n}+1\mod t$ splits when $t$ is a prime satisfying $t=1\mod 2n$.
-In the vectorized case, each ciphertext encrypts a vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$ and the homomorphic operations are single-instruction multiple-data (SIMD) instructions.
+Vectorized SHE.
+When a plaintext modulus $t$ is a prime satisfying $t=1\mod 2n$, these SHE schemes support operations on plaintext vectors as well as polynomials.
+In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation performed by each component of plaintext vectors.
 Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
 SHE operations and costs.
-BFV’s $\mathsf{Eval}$ calls five subroutines:
+SHE schemes we consider support following operations.
   • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$.
@@
 We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
-Ciphertext rotation and ciphertext multiplication require an evaluation key since the secret changes with each rotation and ciphertext multiplication.
-This operation is called “key-switching” for rotations and “relinearization” for ciphertext multiplications.
-Table 1 has concrete run times and noise growth for BFV.
+In Table 1 we show concrete run times and noise growth for BFV, an efficient SHE scheme.
 In general, addition and plaintext multiplication are efficient and addition has minimal noise growth.
@@
 Therefore, minimizing the number of ciphertext multiplications and ciphertext rotations is crucial for scalable applications.
+Evaluation keys.
+Ciphertext rotation and multiplication require evaluation keys that depend on a secret key.
+Additionally, each rotation requires a separate evaluation key.
+Usually, these keys can be reused, but As mentioned later, in Wally  the client must send fresh keys for each query.
+Therefore, these keys contribute to the client’s request size.
 Operation Time (ms) Noise added (bits)
-$\mathsf{CtCtAdd}$ $070$absent0$\approx$ 0≈ 0
-$\mathsf{PtCtMult}$ $0.09$ $22$
-$\mathsf{CtCtMult}$ $12.1$ $20$
-$\mathsf{CtRotate}$ $3.6$ $\approx 0$
+$\mathsf{CtCtAdd}$ $0.004$ $0.5$
+$\mathsf{PtCtMult}$ $0.02$ $20$
+$\mathsf{CtCtMult}$ $2.5$ $26$
+$\mathsf{CtRotate}$ $0.5$ $0.5$
 Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation.
-The polynomial degree $n$ is $8192$, the ciphertext modulus $Q$ has $150$ bits, and the plaintext modulus $t$ has $20$ bits.
-Time costs are measured with the SEAL library [CLP17] version 3.7.2 on a single core in a AWS t2.2xlarge instance.
-Differential privacy
-We use the standard notion of differential privacy (DP) [DMNS06, DKM+06]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
+The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits).
+$\mathsf{CtRotate}$ noise is measured on repeated rotations and the others are from two input ciphertexts with the same noise levels.
+Time costs are measured with the swift-homomorphic-encryption library[7] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
+Differential privacy.
+We use the standard notion of differential privacy (DP) [18, 17]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
 Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
 Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations.
-We use basic composition in DP [Vad17, Lemma 2.3].
+We use basic composition in DP [40, Lemma 2.3].
 Lemma 1.
@@
 Anonymous networks.
-Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [GRS99], or mixnets [Cha81].
+Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [26], or mixnets [12].
 Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
-Private information retreival.
+Private information retrieval.
 Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation.
 Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database.
-PIR protocols can be for index retrieval or retrieval by keyword [CGN98].
-In this paper, we focus on single-server PIR protocols [KO97] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard.
-We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [ACLS18] and [ALP+21].
-There is a notion of differentially-private PIR in the multi-server setting [TDG16, AIVG22], where the database is stored across non-colluding servers while guaranteeing differential privacy.
+PIR protocols can be for index retrieval or retrieval by keyword [14].
+In this paper, we focus on single-server PIR protocols [33] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard.
+We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [5] and [4].
+There is a notion of differentially-private PIR in the multi-server setting [38, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy.
 This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
-3 System Design
+3 System Overview
 Figure 1: High-level overview of Wally.
-Throughout an epoch, clients sends encrypted real and fake queries at random time slots to an anonymization network which “washes” queries of identifying information.
+Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
 A high-level design of Wally is in Figure 1.
@@
 Wally runs in epochs of time; within each epoch, a fixed number of clients make queries.
+Within each epoch, Wally performs a search that outputs the most relevant results to client queries, similar to insecure search engines.
 Wally does not require any synchronization across clients.
-Each client can decide to participate or skip an epoch.
-However, we require a minimum of clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch.
-Overall, privacy loss is the sum of the loss due to each epoch in which the client has participated.
-The correctness guarantee is that all clients’ queries are processed by the end of the epoch.
-Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a can participate at max $M$ epochs.
-Each epoch is divided into second-long time slots, and privacy is guaranteed by the following:
+Each client can independently decide to participate or skip an epoch.
+However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch.
+Overall, privacy loss for a particular client is the sum of the loss due to each epoch in which the client has participated.
+Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a client can participate at max $M$ epochs.
+The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
+In Wally Each epoch is divided into time slots, say a second long, and privacy is guaranteed by the following:
+  • Hiding the content of the queries (somewhat homomorphic encryption).
   • Hiding the real queries of clients among fake queries (differential privacy).
-  • Hiding the content of real and fake queries (somewhat homomorphic encryption).
-  • Removing the identifiable information from each query and shuffling the batch of queries received within a second (anonymization network).
+  • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
   • Hiding the information revealed by queries order and time of arrival (random query schedule)
-Within each epoch, Wally performs a semantic search that outputs the most relevant results to client queries, similar to insecure search engines.
 Embedding-based search.
-Wally utilizes (vector) embeddings for semantic search.
+Similar to state of the art insecure semantic search and previous secure search, Wally utilizes (vector) embeddings for semantic search.
 Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats.
@@
 Therefore, Wally can support all these data types.
-Wally is compatible with all the recent embedding models, such as BERT [DCLT19].
-In the following sections, we introduce the insecure version of Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
+Wally is compatible with all the recent embedding models, such as BERT [15].
+In the following sections, we introduce the insecure version of semantic search in Wally.
+We then highlight the privacy issues in this version.
+Finally, we discuss how we overcome these privacy issues.
 3.1 Insecure Search Pipeline
-Once the client query and the server index are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query.
+In the offline phase, the server maps each document to the corresponding embedding.
+At the time of a query the client embeds the query into an embedding.
+Once the client query and the server documents are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query.
 Our two-step approximate nearest neighbor protocol has the following flow:
   • Server initialization.
-At initialization, the server divides the entries into $K$ clusters using $K$-means clustering.
-A heuristic-based technique partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$ and outputs these clusters with their centroid embeddings $c_{1},\cdots,c_{K}$.
+At initialization, the server divides the documents (using their embeddings) into $K$ clusters using K-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$.
 The server sends the centroid embeddings to the client.
   • Pick nearest centroid.
-To make a query, the client first finds the cluster nearest to the query embedding.
-For this, the client computes a cosine similarity between the query and each centroid embedding and picks a cluster whose centroid has high cosine similarity.
+To make a query, the client first locally finds the cluster nearest to the query embedding.
+For this, the client computes a similarity between the query and each of the centroid embeddings and picks a cluster whose centroid has high similarity.
   • Server computation.
-The client then sends the nearest cluster ID $C^{\prime}$ and query embedding $q^{\prime}$ to the server.
-The server computes cosine similarity between the query embedding and each embedding in cluster $C^{\prime}$.
+The client then sends the nearest cluster index $C^{\prime}$ and query embedding $q^{\prime}$ to the server.
+The server computes similarity between the query embedding and each embedding in cluster $C^{\prime}$.
 The server returns the similarity scores to the client.
@@
 Therefore, finding a balance is important for the overall usability of the system.
-The above protocol is not private. To ensure that the protocol does not reveal any information about the client, it must hide the following information from the server:
+The above protocol is not private. To ensure that the protocol does not reveal any information about the client’s query, it must hide the following information from the server:
 […]
@@
 For example, if the client query is an image, the cluster ID might reveal whether the image is a landmark, e.g., a castle or a bridge.
-  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about client queries.
+  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
 We show how we hide the above data below and in Section 6.
-Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
-3.2 Hiding the Query Embedding with SHE
-To hide a query embedding, we rely on vectorized, RLWE-based SHE introduced in Section 2.
-The client query is an encrypted vector, and the server database is encoded as an array of plaintext vectors.
-The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
+3.2 Hiding the Query Embedding
+We rely on efficient vectorized, RLWE-based SHE described in Section 2 to hide a query embedding.
+At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster.
+The server then returns encrypted scores to the client.
+SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient.
+Overall, in our implementation, the client query and server response are a single RLWE ciphertext of sizes 226kB, and 53kB, respectively, (the former with evaluation keys) and it only takes 19 ms for the server to compute similarity scores for a single cluster homomorphically.
+3.3 Hiding the nearest cluster
+By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster.
+Hiding the nearest cluster is a significant performance bottleneck in previous schemes like Tiptoe.
+Specifically, to conceal the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database (all the clusters).
+Otherwise, skipping clusters will reveal to the server that the client is not interested in them.
+Computing over the entire database hides the nearest cluster in the most robust possible sense, but this results in a prohibitively high server computation cost.
+Instead, we propose an approach based on differential privacy.
+The key advantage is that, in expectation, the server computes encrypted similarity scores only against a few clusters per query.
+Our approach results in a significant improvement in server computation.
+An additional advantage of our approach is a reduction in request size; concretely, the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe.
+Conversely, differential privacy is a weaker privacy notion than the full, cryptographic obliviousness Tiptoe achieves.
+The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making the queries at any given time.
+Therefore, we propose a novel mechanism to hide a particular client query among the crowd of queries from all the clients.
+Our proposed approach provides $(\epsilon,\delta)$-differential privacy.
+This approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability.
+In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
+As the first step, we require all the traffic (queries and responses) to and from the server to route through the anonymization network.
+Intuitively, one would expect that doing this would hide the query in the crowd of queries, as queries do not carry identifiable information.
+However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can learn information about a client.
+For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
+Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client.
+For example, the queries towards each epoch’s start could belong to a particular user.
+To achieve the desired privacy, we make the following two additional changes, these changes are highlighted in Figure 3:
+  • Fake queries.
+In addition to real queries, each client makes a few fake queries.
+Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated.
+In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries.
+For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks one of the clusters.
+  • Random query schedule.
+The client makes queries at random instances within the epoch.
+For this, we assume that each epoch is divided into slots, each one second long.
+For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
+Intuitively, these two changes suppress the leakage, as mentioned above.
+In Section 6, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
+3.4 Hiding Metadata Index
+After learning the index of the nearest entry, the client retrieves metadata associated with it.
+When metadata per entry is small, the server returns metadata of the entire cluster with each query.
+Note that this leaks no extra information because the server already learns the cluster ID for each query.
+The client then locally selects the metadata associated with the most relevant entry.
+However, if the metadata is quite large, the client can run Keyword PIR to fetch the metadata of only the most relevant entry within a cluster.
+3.5 Keyword Search
+Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword.
+Using a random hash function, the server divides the database into random $K^{\prime}$ buckets (instead of clusters).
+For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket.
+The differential privacy part remains the same as in semantic search.
+That is, clients make fake queries for random buckets at a random schedule.
+4 Efficient SHE-Based Instantiation
+In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, the an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls.
+In this section, we discuss the details of our efficient SHE-based instantiation.
+4.1 SHE details in semantic search
+The server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
 Server database.
-The server divides the database, of size $N$, into $K$ clusters each with at most $N^{\prime}$ embeddings.
+Recall that the server database is divided into K clusters.
+Each cluster consists of at most $N^{\prime}$ embeddings.
 For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size.
@@
 Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$.
-For now, we assume that each diagonal of a slice is a separate plaintext vector of size $n=d$.
-That is each slice is encoded as $d$ plaintexts vectors and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$.
+For now, we assume that the plaintext dimension size $n=d$ and each diagonal within a slice is a separate plaintext.
+That is each slice is represented as $d$ plaintexts vectors, and $j$’th vector consists of entries $p_{j}=[e_{0}^{j},e_{1}^{j+1},e_{2}^{j+2},\cdots,e_{d-1}^{j+d-1}]$.
 This means each plaintext vector consists of a single index from each embedding in a slice in increments of one.
 The client encrypts the query embedding $\hat{q}=[q^{0},q^{1},\cdots,q^{d-1}]$ in a single SHE ciphertext.
-Dot product computation.
-Recall, from Section 2, for cosine similarity, the server needs to compute a dot-product between two embeddings.
+Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
+SHE dot-product computation.
+Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them.
 To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$.
@@
 The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice.
-Figure 2 shows an example server computation for a single slice for $d=4$.
-This diagonal method is to minimize homomorphic rotations [HS14, HS18].
+Figure 2 shows an example server computation for a single slice for $n=d=4$.
 Database and query packing.
-In the above description, we assume that the polynomial dimension $n$ equals embedding dimension $d$.
+The above description assumes that the polynomial dimension $n$ equals embedding dimension $d$.
 However, $n$ is often significantly larger than $d$.
@@
 Our scheme packs diagonals across the slices in a single plaintext vector.
-Concretely, the $j$-th plaintext vector in a cube holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices.
+Concretely, a cube’s $j$-th plaintext vector holds $(j\mod d)$-th diagonals of $\lfloor{n/d}\rfloor$ slices.
 In all the practical databases that we have considered, the number of slices per cube $N^{\prime}/d$ is smaller than $\lfloor{n/d}\rfloor$.
@@
 Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
-Recall from Section 2 that homomorphic rotation is a costly operation.
-Even after packing, the server still performs $d$ rotations.
-Looking ahead to Section 6, we explain how we reduce the number of rotations to $2\sqrt{d}$ by using the standard the baby-step giant-step optimization [HS18].
+Recall from Section 2 that homomorphic rotation is costly.
+In the above explanation, the server performs $d$ rotations even after packing to align the query.
+In Wally  we use the standard baby-step giant-step (BSGS) optimization [27] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step).
+See Algorithm 1 in [27] for the BSGS algorithm.
 Float-to-integer encoding.
 Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats.
-Wally converts each float into an integer by multiplying with the appropriate scaling factor and only taking the integer part of the result.
+In Wally we convert each float into an integer by multiplying it with the appropriate scaling factor having the plaintext modulus be large enough to perform the computation over the integers.
 The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
 Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
-3.3 Hiding the nearest cluster
-By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster.
-Hiding nearest cluster is a significant performance bottleneck in state-of-the-art schemes like Tiptoe.
-Specifically, to hide the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database.
-Otherwise, skipping clusters will reveal to the server that the client is not interested in them.
-Computing over the entire database hides the nearest cluster in the strongest possible sense, but this results in a prohibitively high server computation.
-Instead, we propose an approach based on differential privacy.
-The key advantage is that, in expectation, the server computes encrypted similarity scores against a few clusters per query.
-Our approach results in a significant improvement in server computation.
-An additional advantage of our approach is that the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe.
-Conversely, differential privacy is a weaker privacy notion than the full, cryptographic, obliviousness Tiptoe achieves.
-The starting point of our solution is that in large-scale systems, sufficiently many clients are making the queries at any given time.
-Therefore, we propose a novel mechanism allowing clients to hide their queries among the crowd of all queries.
-Our proposed approach is based on $(\epsilon,\delta)$-differential privacy.
-The approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability.
-In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
-Intuitively, one would expect that routing all the traffic through an anonymization network would provide sufficient privacy as queries do not carry identifiable information.
-However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can potentially learn information about a client.
-For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known to not access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
-Our protocol randomizes queries to guarantee differential privacy. Specifically, we make the following two changes:
-  • Fake queries.
-In addition to real queries, each client makes a few fake queries.
-Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated.
-In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries.
-For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks on of the clusters.
-  • Random query schedule.
-The client makes queries at random instances within the epoch.
-For this, we assume that each epoch is divided into a slots, each one second long.
-For each query (real or fake), the client independently picks a random slot and only makes the query at that slot.
-In Section 5, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
-3.4 Hiding Metadata Index
-After learning the index of the nearest entry, the client might want to retrieve metadata associated with it.
-When each entry has a small amount of metadata, the server returns metadata of the entire cluster with each query.
-Note that this leaks no extra information because the server already learns the cluster ID for each query.
-The client then locally picks the metadata associated with the most relevant entry.
-However, if the metadata is quite large, the client can run PIR to fetch the metadata of only the most relevant entry.
-3.5 Keyword Search
-Wally also supports keyword search in which the client is interested in retrieving data associated with a keyword.
-Using a hash function, the server database is divided into random $K^{\prime}$ buckets (instead of clusters).
-For each query, the client locally finds the bucket in which the keyword falls and then uses the keyword PIR scheme to fetch a specific record within a bucket.
-Wally uses a keyword PIR construction based on cuckoo hashing and index PIR.
-At a high level, at initialization, the server uses the cuckoo hashing to map the $N$ database elements into a table of size $O(N)$ using two random hash functions.
-The guarantee is that if a keyword is present in the database, it must be at index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table.
+4.2 SHE details in keyword search
+The server database consists of $N$ (keyword, value) entries.
+Recall that the server divides these entries into $B$ buckets using a random hash function.
+Within each bucket, in Wally  we use a keyword PIR construction based on cuckoo hashing and index PIR.
+At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions.
+The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table.
 The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data.
-Wally uses MulPIR [ALP+21] as the underlying index PIR, an efficient single-server PIR scheme based on SHE.
+Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE.
 We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
-Each bucket consisting of $N/K^{\prime}$ entries represented as $\sqrt{N/K^{\prime}}\times\sqrt{N/K^{\prime}}$ matrix $\mathbf{D}$.
+Each server bucket consisting of $N/B$ entries represented as $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$.
 MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial.
@@
   1. 1.
-    First expand $\hat{q}$ to two $\sqrt{N/K^{\prime}}$-length ciphertext vectors.
+    First expand $\hat{q}$ to two $\sqrt{N/B}$-length ciphertext vectors.
 Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$.
-To expand, the server uses the oblivious expansion algorithm given in [ALP+21].
-This algorithm requires $2\sqrt{N/K^{\prime}}$ homomorphic substitutions.
+The server uses the oblivious expansion algorithm given in [4] to expand.
+This algorithm requires $2\sqrt{N/B}$ homomorphic substitutions.
   2. 2.
-    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/K^{\prime}$ plaintext-ciphertext multiplications and $N/K^{\prime}-\sqrt{N/K^{\prime}}$ ciphertext additions.
+    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/B$ plaintext-ciphertext multiplications and $N/B-\sqrt{N/B}$ ciphertext additions.
   3. 3.
     Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response.
-This step uses $\sqrt{N/K^{\prime}}$ ciphertext-ciphertext multiplications and $\sqrt{N/K^{\prime}}-1$ ciphertext additions.
+This step uses $\sqrt{N/B}$ ciphertext-ciphertext multiplications and $\sqrt{N/B}-1$ ciphertext additions.
 to compute.
-The differential privacy part remains the same as in semantic search.
-That is, clients make fake queries for random buckets at a random schedule.
-Section 6 discusses various optimizations we implemented for MulPIR.
-Overall, our optimizations result in around a three times reduction in server computation and a two-three times reduction in request size.
-[…]
-Figure 4: Simplified protocol for Wally.
-Each client generates queries with Algorithm 2 and sends them to the anonymization network, $\mathcal{AN}$, according to the randomized schedule described in Algorithm 4.
-Cluster $j$ receives anonymous queries denoted as $(\tilde{q}^{j}_{i})$ and computes each response $(r^{j}_{i})$ via SHE.
+Section B discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
 Algorithm 1 $\mathsf{ServerInit}$
@@
 5 Wally Protocol
-In this section, we describe the complete Wally protocol for a single epoch with $U$ clients.
+In this section, we describe the complete Wally protocol for a single epoch in semantic search.
+We assume that $U$ honest clients are present to query within that epoch.
+Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section B.
 Server initialization.
 The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts.
-Specifically, the algorithm first divides the input embeddings into $K$ clusters using K-means clustering.
-We pick $K$ such that the maximum cluster size is less than SHE polynomial dimension $n$ with high probability, which ensures that for each cluster all the embeddings can be encoded into $d$ plaintext vectors.
+Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering.
 The algorithm then iterates over each cluster separately.
-Recall each cluster consists of $|C[i]|/d$ column-wise $d*d$ matrices.
+Each cluster consists of $|C[i]|/d$ column-wise $d*d$ slices.
 The algorithm then fills $d$ plaintext vectors for each cluster.
 In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices.
-Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts and rotating each group in increasing multiples of $-g$.
-Note that these rotations are performed on plaintexts so their cost are negligible.
+Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$.
+These rotations are required for the baby-step giant step (BSGS) optimization [27].
+Note that these rotations are performed on plaintexts, so their cost are negligible.
 Query generation.
-The client uses Algorithm 2 to generate queries and their schedule at the start of an epoch.
-The client locally stores the set of cluster centroids $c$ that the server generates during initialization.
+Each client uses Algorithm 2 to generate queries and their schedule at the start of an epoch.
+The client locally has a set of cluster centroids $c$ that the server generates during initialization.
 The algorithm first generates independent queries for at most $\Delta$ real query embeddings.
@@
 This is done to take advantage of SHE packing.
-The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
+The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and rotation evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
 Next, the algorithm generates fake queries.
@@
 To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$.
+Note that the client must generate a fresh evaluation key to keep each query anonymous.
+Also, each evaluation key must include a rotation key for steps $1$ and $g$, because the server computation involves rotating by these steps.
 The algorithm, then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries.
@@
 During each epoch, the server receives multiple queries for each cluster.
-Towards the end of epoch, the server’s view can be considered as a noisy histogram over all clusters.
+Towards the end of the epoch, the server’s view can be considered as a noisy histogram over all clusters.
 We consider this noisy histogram as an output of the DP mechanism.
@@
 Now consider the cluster $b$; the two databases differ by $\Delta$ real queries for this cluster. By the following, lemma $M_{b}$ provides $(\epsilon,\delta)$ privacy on cluster $b$.
-Lemma 2 (Theorem 13 [GKMP20]).
+Lemma 2 (Theorem 13 [25]).
 […]
@@
 $\sum_{j}\epsilon_{j}=\frac{\epsilon}{\Delta}\sum_{j}\Delta_{j}\leq 2\epsilon.$
-Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ.
-We emphasize that the proof is over a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
+Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ. ∎
+We emphasize that the theorem is stated for a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
 6.2 Generating view of the server
@@
 After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
-6 Optimizations
-Here we present our optimizations for semantic and keyword search.
-Our optimizations revolve around the [Bra12, FV12] and BGV [BGV14] SHE schemes.
-Further, we describe several optimizations to the MulPIR scheme [ALP+21].
-First, we describe the standard RNS optimization in SHE [GHS12] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [HLHD22].
+7 Implementation and Performance Evaluation
+In this section, we demonstrate Wally’s concrete performance via a set of experiments.
+We also showcase Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
+7.1 Implementation
+We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [7].
+Following the library, we use BFV [21] to implement the SHE parts Wally.
+However, our protocol could easily be implemented using other SHE schemes such as BGV [11].
+BFV parameters
+The choice of BFV parameters provides a trade-off between computation and communication in Wally.
+At a high level, larger parameters provide better computation but result in high request size.
+We tried multiple sets of parameters and picked the one that provided the best performance.
+We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits.
+These parameters allow a precision of six bits.
+To increase the precision, we use plaintext CRT optimization described in Section B.
+Recall that this optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query).
+For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$.
+This gives us a precision of 13 bits.
+Differetnial privacy parameters
+Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries.
+For a fixed failure probability $\delta$, smaller $\epsilon$ provides stronger privacy but worse performance because each client is required to make more fake queries.
+We set $\delta=2^{-20}$, which means DP will fail only once a million and $\epsilon=1$, providing a decent privacy guarantee.
+7.2 Performance Evaluation
+Experimental setup.
+We ran server-side computations on an Intel Xeon w3-2423 instance with 32GB of RAM and 6 cores.
+End-to-end performance.
+Metric Low prec. Mid prec. Tiptoe
+Query Size (MB) 0.67 0.84 16.2
+Response Size
+(MB) 0.16 0.18 3.2
+Time to process
+queries 39 mins 44 mins 2.8 years
+Precision (bits) 6 13 4
+Privacy DP DP SHE
+Table 2: Performance comparison of Wally and Tiptoe for database of 35 millions entries.
+Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query).
+Throughput for Wally is measured for 8 million real queries, or 24 million in total.
+All Wally responses are computed as the encrypted scores plus $44$ KB of metadata.
+Table 2 describes our main experiments compared to Tiptoe.
+The server database consists of 35 million entries divided into around 8500 clusters.
+We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$.
+We assume each client makes $\Delta=2$ queries per epoch for these experiments.
+Moreover, there are four million honest clients present in an epoch.
+This results in eight million real queries in an epoch.
+For Wally, we calculated the overhead by using Theorem 2. The fake query overhead is derived by the expected number of fake queries for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
+$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
+where $K$ is the total number of clusters and $U$ is the total number of clients.
+For our selected parameters, for each real query, the client is required to make just under two fake queries.
+Therefore, the total number of queries within an epoch is 24 million.
+In Table 2, we included fake query overhead in query request, response and total computation in Wally .
+Epoch size in Wally equals the total time it takes to process all the queries (real and fake).
+For Tiptoe, we assume there are eight million queries the server processes one by one.
+Tiptoe requires running experiments on real compute nodes; due to resource limitations, we carefully estimate their results using Table 7 in [29].
+Before discussing Wally’s performance, we emphasize that Wally requires the following two extra assumptions not required in Tiptoe.
+  • Wally requires sufficiently many honest clients during an epoch.
+  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
+In other words, Wally trades off scalability for privacy.
+Wally has significantly better performance than Tiptoe across all the metrics.
+Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe.
+This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters.
+Similarly, the response size is 3.1-3.6 smaller.
+Wallyś efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe.
+For example, to process eight million queries, the server with 320 cores requires around 39 minutes in Wally .
+In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
+Scaling DP with $\epsilon$.
+Figure 4 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients.
+This is derived by dividing Equation 1 by the number of clusters $K$.
+Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
+Figure 4: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
+Precision.
+Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section B.
+The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$.
+Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
+Precision (bits) 6 13 29
+Query (KB) 226 283 901
+Response (KB) 53 63 196
+Server comp. (ms) 19 35 110
+Table 3: Here we list the query and response size per each query (real and fake).
+Each response size includes $44$kB of metadata.
+Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
+Scaling with the number of clients.
+Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system.
+We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients.
+As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads.
+Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
+Number of Clients .5M 1M 2M 10M
+Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
+Exp. MB ($\Delta=1$) $1.9$ $1.08$ $0.65$ $3.11$
+Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
+Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$ $0.61$
+Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
+7.3 Keyword Query Performance
+Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [7].
+Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4].
+Each response size was 27KB and the query sizes were 368 KB, 539 KB, 539 KB, and 624 KB for each respective row in the table.
+Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
+Database entries Entry Size (Bytes) Time (ms)
+100k 10 166
+100k 100 465
+1 million 10 565
+1 million 100 1700
+Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section B.
+8 Conclusion
+Wally demonstrates that private search systems can be scalable to systems with many users.
+Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters.
+Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases.
+This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.
+9 Open science and ethics discussion
+9.1 Open Science Policy
+We have ensured open science policies by taking the following steps: We develop our system using an open-source library.
+The paper discusses and provides all the research techniques we used to develop the end system.
+We provide open access to the paper by sharing it on Arxiv Once accepted, we will reveal part of the code that will allow reproducibility.
+The entire protocol is a simple extension of the code we will open source, but we cannot provide a full implementation of the end-to-end deployed system due to our company’s IP policy.
+9.2 Ethics
+This paper complies with all the ethics guidelines.
+Specifically, we do not use private data or code without developers’ permission or run experiments with people.
+Our protocol protects user privacy in search systems without affecting the system’s usability; therefore, it benefits all parties.
+We also took essential disclosures from our employer to share the results given in our paper.
+Appendix A Related Work
+Protocols like Tor [16] and iCloud Private Relay [6] do not provide query privacy.
+Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
+Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness.
+Tiptoe [29], the most recent and performant, has been discussed above in-detail.
+Other, less efficient solutions, include Coeus [1] as well as Hers [20], with the latter providing privacy for the server’s database as well.
+Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[8, 35], and including a client public key in each query [23].
+Another related area of work is differentially-private multi-server PIR schemes [38, 2].
+In both works, servers learn differentially private, noised histograms over queries.
+The former, [38], has clients generate fake queries whereas the latter, [2], has the servers inject fake queries.
+Both require a secure multiparty computation over the entire database.
+Notably, [2] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation.
+Their offline and online computation grows linearly with the size of the server’s database.
+Differential privacy has also been applied to multiple anonymous messaging systems [41, 39, 34].
+These are large (many server) mix-net-like protocols where the servers add in fake queries to hide user messages with a differential privacy guarantee.
+Though they are for a different application, their use of differential privacy is similar to Wally’s.
+Appendix B Optimizations
+Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR).
+Our optimizations revolve around the [10, 21] and BGV [11] SHE schemes.
+Further, we describe several optimizations to the MulPIR scheme [4].
+First, we describe the standard RNS optimization in SHE [24] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [31].
 RNS form.
@@
 Dropping ciphertext LSBs.
-We use Cheetah’s [HLHD22] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext.
+We use Cheetah’s [31] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext.
 This technique compresses the response size further than simply modulus-switching down, but it also adds noise.
-Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [HHC+23]), but we derive the proper analysis as follows.
+Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [30]), but we derive the proper analysis as follows.
 Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$.
@@
 Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined our BFV plaintext CRT optimization described in the next subsection.
-6.1 Approximate Nearest Neighbors Optimizations
+B.1 Approximate Nearest Neighbors Optimizations
 The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size.
@@
 Minimal rotation keys and bandwidth.
-Wally uses the “baby-step giant-step” (BSGS) optimization [HS18] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant).
-See Algorithm 1 in [HS18] for the BSGS algorithm.
+Wally uses the “baby-step giant-step” (BSGS) optimization [27] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant).
+See Algorithm 1 in [27] for the BSGS algorithm.
 Additionally, we are able to use two rotation keys since $d<n/2$ in our use case.
 This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$.
-For key-switching, and relinearization in PIR below, we use the hybrid-GHS [GHS12] key-switching strategy described in-detail by Kim et al.
-[KPZ21].
+For key-switching, and relinearization in PIR below, we use the hybrid-GHS [24] key-switching strategy described in-detail by Kim et al.
+[32].
 Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision.
-The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [APS15] [^4] https://github.com/malb/lattice-estimator.
+The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [3] [^3] https://github.com/malb/lattice-estimator.
 A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
 Re-using BFV evaluation keys via the plaintext CRT.
-Wally achieves minimal bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space.
-The CRT on plaintexts allows us to keep a set amount of plaintext bits while re-using one set of evaluation keys across all plaintext spaces since BFV’s evaluation keys are independent of the ciphertext modulus, unlike BGV evaluation keys.
-Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{i}$ for NTT-friendly primes $\{t_{i}\}$.
-The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}\cdots t_{l}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
-Therefore, we can drop the ring dimension $n^{\prime}=4096$ and use two ciphertexts, modulo $t_{0}=40961\approx 2^{15}$ and modulo $t_{1}=65537\approx 2^{16}$ for $t_{0}t_{1}\approx 2^{31}$, half the bandwidth from ciphertexts, and use evaluation keys in dimension $n^{\prime}$, or one fourth the bandwidth from evaluation keys.
-Concretely, this leaves us with $113$kB for ciphertexts and $170$kB for evaluation keys, $283$kB in total, over a 3×3$\times$3 × improvement in query size without sacrificing plaintext precision.
+Wally achieves minimal query bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space.
+The CRT on plaintexts allows increase the plaintext precision without increasing evaluation key size since BFV’s evaluation keys are independent of the plaintext modulus.
+Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{0}$ and $\mathbf{\widetilde{v}}\mod t_{1}$ for NTT-friendly primes $\{t_{0},t_{1}\}$.
+The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
 In general, the plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: six polynomials versus 2 polynomials.
@@
 One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
-6.2 PIR Optimizations
+B.2 PIR Optimizations
 Uneven dimensions.
 In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication.
-For example, SEAL [CLP17] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [MR23].
+For example, SEAL [13] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [37].
 Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$.
@@
 MulPIR is an optimized query expansion which originated with Angel et al.
-[ACLS18].
+[5].
 In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching).
@@
 Lazy rescaling in BFV multiplication.
-We apply the lazy rescaling technique of Kim et al [KPZ21] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$.
+We apply the lazy rescaling technique of Kim et al [32] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$.
 The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$.
-Kim et al [KPZ21, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
+Kim et al [32, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
 Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
-Handling Very Large Entries
+Optimization for Very Large Entries
 In MulPIR the database is represented as a plaintext matrix $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$. Conventionally, the response is computed as
 […]
-This approach results in $n\cdot t$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry.
+This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry.
 The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large.
-But it will blow up with extremely large entries (e.g.
-when they are high-resolution photos).
+But it will blow up with extremely large entries (e.g., when they are high-resolution photos).
 To overcome this challenge, we can swap the order of the computation as
 […]
-The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size and $t$.
-This optimization reduce the computation overhead when $t$ is larger than the first dimension $d_{1}$.
+The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size.
+While the inner products still takes $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications.
+Therefore, this optimization can reduce the computation overhead when the entries size $t$ is larger than the first dimension size $d_{1}$.
 Keyword PIR.
@@
 BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around.
-We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [BEHZ16] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
-We add a conditional, lazy modular reduce to Harvey’s NTT [Har14].
+We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [9] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
+We add a conditional, lazy modular reduce to Harvey’s NTT [28].
 The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size.
@@
 Note, we saw no improvement on primes near $64$-bits.
-7 Performance
-In this section, we demonstrate Wally’s scalability and concrete performance as seen in experiments.
-We also track Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
-Experimental setup.
-We ran server-side computations on an AWS c5.24xlarge instance with 192GB of RAM, single-threaded. We extrapolate throughput from the single-threaded times seen on this AWS instance.
-End-to-end performance.
-Metric Low prec. Mid prec. Tiptoe
-Exp. query (MB) 0.67 0.84 16.2
-Exp. response (MB) 0.16 0.18 3.2
-Queries/s 2998 1499 47
-8M queries (hrs) 0.4 1.5 47.5
-Fake query 3×3$\times$3 × 3×3$\times$3 × --
-overhead
-Precision (bits) 6 13 4
-Privacy DP DP SHE
-Network req. AN AN --
-Table 2: Wally performance for a parameter set with BFV dimension $n=4096$ on a 35M entry dataset with the maximum cluster size as $4096$ compared to Tiptoe.
-Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query).
-Throughput for Wally is measured for 8M real queries, or 24M in total.
-All Wally responses are computed as the encrypted scores plus $44$kB of metadata.
-Table 2 describes our main experiments in comparison with Tiptoe.
-Here, we use two parameter sets for Wally: first, we use a small BFV parameter set ($n=4096$, $\log_{2}t=15$, and $\log_{2}q=83$) and second, a plaintext CRT (Section 6) parameter set which includes the former.
-The plaintext precision is $(\log_{2}t)/2-(\log_{2}d)/4$ since the inner-product computation grows heuristically as $\sqrt{d}\|\mathbf{\tilde{e}}\|_{\infty}^{2}$ where $\mathbf{\tilde{e}}=\alpha\cdot\mathbf{e}$ is a scaled (unit) embedding vector, $\|\mathbf{e}\|_{2}=1$.
-The former parameter set enables six bits of plaintext precision whereas the latter enables 13 bits.
-We set $\Delta=2$ messages per client per epoch and calculated the DP overhead by using $(\epsilon,\delta)=(1,2^{-20})$ after the loss in Theorem 2 with four million clients per epoch.
-The fake query overhead is derived by the expected number of fake queries, for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
-$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
-where $K$ is the total number of clusters and $U$ is the total number of clients.
-Wally’s ability to cluster the dataset and keep server computation time proportional to the maximum size of a cluster is demonstrated in its small per-query server computation time, $95$ms for each real or fake query, and its relatively small expected query size, $679$kB.
-We keep each cluster size less than the total number of plaintext slots in each BFV ciphertext, $n=4096$.
-Otherwise, Wally can compute similarity scores in parallel if the cluster size is larger than $n$.
-Large meta-data sizes can be retrieved by adding another round of PIR after the similarity computation.
-The throughput times represent how many queries per second can be processed on $K=8545$ clusters during an epoch needed to serve $24$M total queries, real and fake.
-This throughput calculation leads to a minimum epoch size of $44$ minutes serving roughly 3k queries per second.
-This compares to the $47.5$ hours Tiptoe would take to serve $8$M queries.
-Comparing to Tiptoe.
-Wally has the following two extra assumptions compared to Tiptoe:
-  • Wally requires sufficiently many honest clients in the system during an epoch.
-  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short amount of time, e.g., one second time slots.
-In other words, Wally trades off scalability for privacy.
-Such a tradeoff is needed since oblivious, cryptographically secure systems require the server to compute over the entire database and cannot scale to many clients.
-Tiptoe ran its encrypted similarity computation on a cluster of $80$ r5.xlarge AWS instances, each with 4 vCPUs and 32 GB of DRAM, and saw a latency of $11.2$ seconds per query, of which is a significant amount of per-query preprocessing (8.7 seconds).
-Each server held $8$GB to $12$GB of the plaintext database.
-We emphasize that each Tiptoe query computes of the entire database (all clusters), and we estimate their total single threaded time as 80×80$\times$80 × their per query time, or $183$ seconds per query.
-We obtained this estimate by dividing the pre-processing (token) times and the download bandwidths in Table 7 in [HDCZ23] by six since our embedding dimension is 3×3$\times$3 × higher and Tiptoe generates two pre-processing elements per query, “tokens”: one each for a similarity computation and PIR.
-The non-pre-processing times and bandwidths, as well as the token upload bandwidth, are divided by 3×3$\times$3 × again due to the difference in embedding dimension and that they use the same token upload for both the similarity computation and PIR.
-Next, the server compute times are multiplied by $80$ since they use $80$ clusters, yielding an estimated $163$ seconds, single-threaded, for a homomorphic similarity computation, i.e., a Wally query.
-Given this time, we divide it by the number of clusters we expect in Wally, $8545$, to estimate throughput in Table 2.
-In comparison, Wally can process each query in parallel since queries, real and fake, have their computation restricted to a single cluster.
-If there are $U$ clients making at most $\Delta$ queries, then Wally can set a epoch time as the minimal amount of time needed to compute $\phi\times U\times\Delta$ queries.
-Since each client sends a query at a random time over an epoch, the average response time is half the epoch whereas the worst case time is the epoch length.
-We derived Wally’s throughput numbers in Table 2 using $95$ms per query, $U=4$M, $\Delta=2$, and $\phi=2$.
-This leads to $24$M queries in 44 minutes, of which $8$M are real queries, on $K=8545$ clusters.
-Scaling DP with $\epsilon$.
-Figure 5 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients.
-This is derived by dividing Equation 1 by the number of clusters $K$.
-Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
-Figure 5: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
-Precision.
-Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section 6.
-The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$.
-Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
-Precision (bits) 6 13 29
-Query (KB) 226 283 901
-Response (KB) 53 63 196
-Server comp. (ms) 95 190 580
-Table 3: Here we list the query and response size per each query (real and fake).
-Each response size includes $44$kB of metadata.
-Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
-Scaling with the number of clients.
-Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system.
-We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients.
-As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads.
-Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
-Number of Clients 0.5M 1M 2M 10M
-Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
-Exp. MB ($\Delta=1$) $1.9$ $1.08$65.65.65.65 $3.11$
-Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
-Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$61.61.61.61
-Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
-7.1 Keyword Query Performance
-Table 5 shows single-threaded keyword PIR server performance, measured on an M1 Max laptop.
-Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
-The request size and the response size for each entry in the table is 728KB and 39KB, respectively.
-Database Entries Entry Size (bytes) Time (s)
-100K 10 0.2
-100K 100 0.5
-1M 10 0.7
-1M 100 2
-10M 10 2.7
-100M 100 6.3
-Table 5: Keyword-PIR, server run times for Wally. These times include the optimizations in Section 6.
-8 Conclusion
-Wally demonstrates that private search systems can be scalable to systems with many users.
-Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters.
-Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases.
-This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.

### v4

@@
-Wally: An Efficient Private Search Engine
+Scalable Private Search with Wally
 Abstract
@@
 As a result, performance degraded proportionally with the number of entries in the database.
-In Wally, we get rid of this limitation.
-Specifically, for each query the server performs cryptographic operations only against a few database entries.
+In Wally, we remove this limitation.
+Specifically, for each query the server performs cryptographic operations against only a few database entries.
 We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants.
@@
 Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases.
-Concretely, Wally can process eight million queries in just 39 mins.
-That is around four orders of magnitude less than the state of the art.
+Concretely, Wally can process eight million queries in 117 minutes, or just under two hours.
+That is around four orders of magnitude faster than the state of the art.
 1 Introduction
@@
 However, differential privacy is an accepted standard for strong privacy [16].
-Specifically, every client in Wally is guaranteed that all the statistical inferences are insensitive to their data.
+Specifically, every client in Wally is guaranteed that any statistical inferences are insensitive to their data.
 In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
@@
 The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch.
-Server side, the database is divided into $K$ clusters.
+The server divides the database into $K$ clusters.
 Each query consists of a relevant cluster index, and the server response is all the entries in the cluster.
@@
 Even though queries are anonymized, they do not provide query privacy.
-The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (if queries arrive earlier than the others).
+The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries).
 To prevent this leakage, Wally implements the following two changes at the client.
@@
 This reduces the response because the size of encrypted scores is significantly smaller than the entries.
-For a keyword search, the client includes encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword.
+For a keyword search, the client sends an encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword.
 To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data.
 We propose various optimizations that significantly improve performance for both kinds of searches.
-Overall, a database with 35 million entries Wally can process eight million real queries in just 39 minutes on 320 cores set up, a drastic improvement over the state-of-the-art, which would have required years to process the same queries.
-Additionally, the size of each request is 53 times smaller and the size of each responses is 3.5 times smaller compared to Tiptoe.
+Overall, for a database with 35 million entries Wally can process eight million real queries in 117 minutes on 320 cores, a drastic improvement over the state-of-the-art, which would have required years to process the same queries.
+Additionally, the size of each request is 53 times smaller and the size of each response is 3.5 times smaller compared to Tiptoe.
 Our contributions are as follows:
   • We describe Wally’s design, which enables privacy-preserving large-scale search engines.
-Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt their queries.
+Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt its queries.
   • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee.
@@
   • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
-Due to space limitation, we discussed related works in Appendix A.
 Limitation.
@@
 Definition 1.
-A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ iid variables.
+A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d.
+variables.
 That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
 2.1 (Somewhat) Homomorphic Encryption
-Fully homomorphic encryption (FHE) is a special scheme allowing arbitrary computation over ciphertexts.
+Fully homomorphic encryption (FHE) is a special cryptosystem allowing arbitrary computation over ciphertexts.
 FHE for arbitrary computation is still very expensive.
-To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports computation up to a fixed depth.
+To achieve practical performance, somewhat homomorphic encryption (SHE, also called leveled FHE) is often used, which only supports a limited number of computations.
 We focus on SHE schemes based on the Ring Learning with Errors (RLWE) problem.
@@
 Here $a$ is picked uniform randomly and $e$ is a noise polynomial with coefficients sampled from a bounded Gaussian distribution.
-A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$ This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$.
+A scheme satisfies its decryption formula: $c_{0}+c_{1}s\mod Q=m+e$.
+This noise polynomial’s coefficients grow as we compute more homomorphic operations on the ciphertext, and we can only decrypt correctly if $\|e\|_{\infty}<Q/2t$.
 We call $\log_{2}(Q)-\log_{2}(2t)$ the parameter’s noise budget, measured in bits.
@@
 When a plaintext modulus $t$ is a prime satisfying $t=1\mod 2n$, these SHE schemes support operations on plaintext vectors as well as polynomials.
-In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation performed by each component of plaintext vectors.
+In the vectorized case, each ciphertext encrypts a plaintext vector $\mathbf{v}\in\mathbb{Z}_{t}^{n}$, and the homomorphic operations are single-instruction multiple-data (SIMD) instructions, that is, the same operation is performed on each component of the plaintext vectors.
 Vectorized SHE is crucial for applications requiring homomorphic linear algebra.
 SHE operations and costs.
-SHE schemes we consider support following operations.
-  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$.
-  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$
-  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$.
-  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$ if $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$ and $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v}^{\prime})$.
+SHE schemes we consider support the following operations. Given $\mathsf{ct}=\mathsf{Enc}(\mathbf{v})$, $\mathsf{ct}^{\prime}=\mathsf{Enc}(\mathbf{v})$,
+  • $\mathsf{PtCtAdd}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
+  • $\mathsf{CtCtAdd}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}+\mathbf{v}^{\prime}$.
+  • $\mathsf{PtCtMult}(\mathsf{ct},\mathbf{v}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
+  • $\mathsf{CtCtMult}(\mathsf{ct},\mathsf{ct}^{\prime})$ returns a ciphertext encrypting $\mathbf{v}\odot\mathbf{v}^{\prime}$.
   • $\mathsf{CtRotate}(\mathsf{ct},r)$ for $r\in[0,n/2)$ returns a ciphertext encrypting
@@
 Differential privacy.
-We use the standard notion of differential privacy (DP) [18, 17]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private, if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
+We use the standard notion of differential privacy (DP) [15, 14]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
 Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
 Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations.
-We use basic composition in DP [40, Lemma 2.3].
+We use basic composition in DP [36, Lemma 2.3].
 Lemma 1.
@@
 Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
-A high-level design of Wally is in Figure 1.
+Figure 1 shows a high-level design of Wally.
 It consists of three entities: the server, the clients, and the anonymization network.
@@
 However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch.
-Overall, privacy loss for a particular client is the sum of the loss due to each epoch in which the client has participated.
-Wally’s privacy notion also requires that each client will make at most $M$ queries in the system’s lifetime; therefore, a client can participate at max $M$ epochs.
+Overall, the privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated.
+Wally’s privacy notion also requires that each client makes at most $M$ queries in the system’s lifetime; therefore, a client can participate at most $M$ epochs.
 The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
-In Wally Each epoch is divided into time slots, say a second long, and privacy is guaranteed by the following:
+In Wally each epoch is divided into time slots, each a second long, e.g., and privacy is guaranteed by the following:
   • Hiding the content of the queries (somewhat homomorphic encryption).
@@
   • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
-  • Hiding the information revealed by queries order and time of arrival (random query schedule)
+  • Hiding the information revealed by queries order and time of arrival (random query schedule).
 Embedding-based search.
@@
 These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space.
-Hence, it is used to find relevant entries in search systems.
+Hence, they are used to find relevant entries in search systems.
 Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication.
@@
 Recall that clustering puts highly correlated documents together in a cluster.
-Therefore, cluster ID reveals the kind of data the client holds.
-For example, if the client query is an image, the cluster ID might reveal whether the image is a landmark, e.g., a castle or a bridge.
+Therefore, the cluster ID can reveal the kind of data the client holds.
+For example, if the client query is a song embedding, the cluster ID might reveal the genre of the song.
   • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
@@
 Our proposed approach provides $(\epsilon,\delta)$-differential privacy.
-This approach guarantees to every client that the maximum privacy loss due to their queries is essentially bounded by $\epsilon$ with all but $\delta$ probability.
+This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $\epsilon$ with all but $\delta$ probability.
 In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
@@
 For example, the queries towards each epoch’s start could belong to a particular user.
-To achieve the desired privacy, we make the following two additional changes, these changes are highlighted in Figure 3:
+To achieve the desired privacy, we make the following two additional changes, highlighted in Figure 3:
   • Fake queries.
@@
 Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword.
-Using a random hash function, the server divides the database into random $K^{\prime}$ buckets (instead of clusters).
+Using a random hash function, the server divides the database into $K^{\prime}$ random buckets (instead of clusters).
 For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket.
@@
 4 Efficient SHE-Based Instantiation
-In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, the an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls.
+In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls.
 In this section, we discuss the details of our efficient SHE-based instantiation.
@@
 Each cluster consists of at most $N^{\prime}$ embeddings.
-For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d*d*d$, recall that $d$ is the embedding size.
-Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d*d$, and each column within a slice is an embedding.
+For simplicity, we assume that for a given cluster $C_{i}$, the dataset is represented as a cube $D_{C_{i}}$ of dimensions $N^{\prime}/d\times d\times d$, recall that $d$ is the embedding size.
+Each cube consists of $N^{\prime}/d$ slices, each of which is a matrix of dimensions $d\times d$, and each column within a slice is an embedding.
 Denote elements of $i$’th embedding as $e_{i}=[e_{i}^{0},\cdots,e_{i}^{d-1}]$.
@@
 Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them.
-To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be like $e_{j}^{i}*q^{i}$.
+To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$.
 The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment.
@@
 Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE.
-We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR schemes
-Each server bucket consisting of $N/B$ entries represented as $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$.
+We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
+Each server bucket consisting of $N/B$ entries is represented as a $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$.
 MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial.
@@
 to compute.
-Section B discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
+Section A discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
 Algorithm 1 $\mathsf{ServerInit}$
 Input: Semantic search database $\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]}$, number of clusters $K$
-Output: Processed semantic $D_{\text{sem}}$,
-Clusters centroids $c$.
+Output: Processed semantic $D_{\text{sem}}$, cluster centroids $c$.
 1:$(\mathbf{c},\mathbf{C})\leftarrow\textsf{K-MeansClustering}(\{(\mathbf{e}_{i},m_{i})\}_{l\in[N]},K)$triangleright dividing semantic data into $K$ clusters
@@
 We assume that $U$ honest clients are present to query within that epoch.
-Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section B.
+Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section A.
 Server initialization.
@@
 The algorithm then iterates over each cluster separately.
-Each cluster consists of $|C[i]|/d$ column-wise $d*d$ slices.
+Each cluster consists of $|C[i]|/d$ column-wise $d\times d$ slices.
 The algorithm then fills $d$ plaintext vectors for each cluster.
@@
 Also, each evaluation key must include a rotation key for steps $1$ and $g$, because the server computation involves rotating by these steps.
-The algorithm, then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries.
+The algorithm then permutes the real and fake queries list $\mathbf{Q}$ and generates a random schedule for the queries.
 To generate the schedule $\mathbf{S}$, for each query $i\in Q$, the algorithm picks a random slot $t$ independently and appends $i$ to $S[t]$.
@@
 Recall that we ensure each cluster is encoded into $d$ plaintexts.
-The algorithm iterates in groups of $g$ plaintexts; within each group, multiply $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts.
+The algorithm iterates in groups of $g$ plaintexts; within each group, multiply the $i$-th plaintext with the $i$-th copy of the rotated query and sum the resulting ciphertexts.
 These iterations yield $h=\sqrt{d}$ ciphertexts.
@@
 We define the curator as Algorithm $M$ that takes input from a database $X$ of requests from $N$ clients and outputs a noisy histogram.
-We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing a client with another.
+We prove the theorem by investigating the affected buckets when we change the input $X$ by replacing one client with another.
 For a particular cluster $b$ in the histogram, we define $M_{b}(\cdot):=R_{b}+\mathcal{NB}(r,p)$, where $R_{b}$ is the total real queries for the cluster $b$ in the database.
@@
 Proof.
-We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard hybrid argument in cryptography.
+We use infinite divisibility of $\mathcal{NB}$ distribution to produce a standard cryptography hybrid argument.
 Hybrid 0: Every client is the same as the real client given in Algorithm 2 except it does not include encrypted embeddings with each request (real or fake).
@@
 Hybrid 1: In this hybrid, we replace all the clients with a single simulator.
-The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ is real queries for $u$-th client.
+The input of a simulator is a list $R=\{R_{u}\}_{u}$, where $R_{u}$ are the real queries for $u$-th client.
 The simulator then simulates each client $u$ with input $R_{u}$ as in Hybrid 0.
-As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$. From the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
+As the simulator internally runs each client $u$ as in Hybrid 0 with input $R_{u}$, from the adversary’s point of view, the distribution of requests is the same as Hybrid 0.
 Hybrid 2: In this Hybrid, the simulator’s input is the same as in the previous Hybrid 1. The simulator is defined as follows.
@@
   2. 2.
-    In Hybrid 1, each simulated client generates a schedule for its queries independently. While in Hybrid 2, the simulator generates a schedule for queries.
+    In Hybrid 1, each simulated client generates a schedule for its queries independently, while in Hybrid 2, the simulator generates a schedule for queries.
 To show that Hybrid 2 and 1 are indistinguishable, we use the definition of infinitely divisible distribution, Definition 1.
-Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)\ \sum_{i=1}^{U}\mathcal{NB}(r/U,p)$.
+Note that negative binomial distribution is infinite divisible because $\mathcal{NB}(r,p)=\sum_{i=1}^{U}\mathcal{NB}(r/U,p)$.
 Thus, the distribution of per cluster fake queries per cluster in Hybrid 2 (Line 2) is the same as in Hybrid 1.
@@
 We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [6].
-Following the library, we use BFV [21] to implement the SHE parts Wally.
+Following the library, we use BFV [18] to implement the SHE parts of Wally.
 However, our protocol could easily be implemented using other SHE schemes such as BGV [10].
@@
 These parameters allow a precision of six bits.
-To increase the precision, we use plaintext CRT optimization described in Section B.
-Recall that this optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query).
+To increase the precision, we use plaintext CRT optimization described in Section A.
+This optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query).
 For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$.
 This gives us a precision of 13 bits.
-Differetnial privacy parameters
+Differential privacy parameters
 Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries.
@@
 Time to process
-queries 39 mins 44 mins 2.8 years
+queries 117 mins 139 mins 2.8 years
 Precision (bits) 6 13 4
@@
 Table 2 describes our main experiments compared to Tiptoe.
+For an end-to-end system, we estimate the network time on a 30Mbps channel.
+This comes out to 74ms for transmitting a query and response over the network for the small precision setting (first column in Tables 2 and 3).
 The server database consists of 35 million entries divided into around 8500 clusters.
@@
 Therefore, the total number of queries within an epoch is 24 million.
-In Table 2, we included fake query overhead in query request, response and total computation in Wally .
-Epoch size in Wally equals the total time it takes to process all the queries (real and fake).
+In Table 2, we included fake query overhead in query request, response and total computation plus estimated network time in Wally.
+The epoch size in Wally is the total time it takes to process all the queries (real and fake).
 For Tiptoe, we assume there are eight million queries the server processes one by one.
@@
   • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
-In other words, Wally trades off scalability for privacy.
+In other words, Wally trades off privacy guarantees for scalability.
 Wally has significantly better performance than Tiptoe across all the metrics.
@@
 Similarly, the response size is 3.1-3.6 smaller.
-Wallyś efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe.
-For example, to process eight million queries, the server with 320 cores requires around 39 minutes in Wally .
+Wally’s efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe.
+For example, to process eight million queries, the server with 320 cores requires around 117 minutes in Wally.
 In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
@@
 Precision.
-Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section B.
+Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section A.
 The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$.
@@
 Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4].
-Each response size was 27KB and the query sizes were 368 KB, 539 KB, 539 KB, and 624 KB for each respective row in the table.
+Each response size was 27KB.
 Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
-Database entries Entry Size (Bytes) Time (ms)
-100k 10 166
-100k 100 465
-1 million 10 565
-1 million 100 1700
-Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section B.
-8 Conclusion
-Wally demonstrates that private search systems can be scalable to systems with many users.
-Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters.
-Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases.
-This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.
-9 Open science and ethics discussion
-9.1 Open Science Policy
-We have ensured open science policies by taking the following steps: We develop our system using an open-source library.
-The paper discusses and provides all the research techniques we used to develop the end system.
-We provide open access to the paper by sharing it on Arxiv Once accepted, we will reveal part of the code that will allow reproducibility.
-The entire protocol is a simple extension of the code we will open source, but we cannot provide a full implementation of the end-to-end deployed system due to our company’s IP policy.
-9.2 Ethics
-This paper complies with all the ethics guidelines.
-Specifically, we do not use private data or code without developers’ permission or run experiments with people.
-Our protocol protects user privacy in search systems without affecting the system’s usability; therefore, it benefits all parties.
-We also took essential disclosures from our employer to share the results given in our paper.
-Appendix A Related Work
-Protocols like Tor [16] and iCloud Private Relay [6] do not provide query privacy.
-Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
+Database entries Entry Size (Bytes) Query size (KB) Time (ms)
+100k 10 368 166
+100k 100 539 465
+1 million 10 539 565
+1 million 100 624 1700
+Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section A.
+8 Related Work
+Anonymization networks alone do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
 Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness.
 Tiptoe [26], the most recent and performant, has been discussed above in-detail.
-Other, less efficient solutions, include Coeus [1] as well as Hers [20], with the latter providing privacy for the server’s database as well.
+Other less efficient solutions include Coeus [1] as well as Hers [17], with the latter providing privacy for the server’s database as well.
 Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[7, 32], and including a client public key in each query [20].
@@
 Though they are for a different application, their use of differential privacy is similar to Wally’s.
-Appendix B Optimizations
+9 Conclusion
+Wally demonstrates that private search systems can be scalable to systems with many users.
+Wally carefully balances somewhat homomorphic encryption together with differential privacy to hide the client’s query and reveal a differentially private histogram over all clients’ traffic to database clusters.
+Further, Wally’s overhead from differential privacy vanishes as the number of users in the system increases.
+This vanishing overhead is crucial to scalability and is potentially useful in many private systems deployable at-scale.
+Appendix A Optimizations
 Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR).
-Our optimizations revolve around the [10, 21] and BGV [11] SHE schemes.
+Our optimizations revolve around the BFV [9, 18] and BGV [10] SHE schemes.
 Further, we describe several optimizations to the MulPIR scheme [4].
@@
 We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$.
-Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined our BFV plaintext CRT optimization described in the next subsection.
-B.1 Approximate Nearest Neighbors Optimizations
+Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
+A.1 Approximate Nearest Neighbors Optimizations
 The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size.
@@
 The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
-In general, the plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: six polynomials versus 2 polynomials.
+In general, plaintext CRT is preferable to increasing the ring dimension since the evaluation key sizes dominate the ciphertext size: 6 polynomials versus 2 polynomials.
 This is true in the bandwidth-optimal key-switching strategy for hybrid GHS key-switching.
@@
 One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
-B.2 PIR Optimizations
+A.2 PIR Optimizations
 Uneven dimensions.
 In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication.
-For example, SEAL [13] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [37].
+For example, Table 1 shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
 Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$.
@@
 The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$.
-Kim et al [32, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
+Kim et al [29, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
 Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
@@
 This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry.
-The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large.
-But it will blow up with extremely large entries (e.g., when they are high-resolution photos).
+The number of ciphertext-ciphertext multiplications grows linearly with $t$ and remains a concretely small value when entries are not large.
+However, it will blow up with extremely large entries (e.g., when they are high-resolution photos).
 To overcome this challenge, we can swap the order of the computation as
 […]
-The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size.
-While the inner products still takes $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications.
-Therefore, this optimization can reduce the computation overhead when the entries size $t$ is larger than the first dimension size $d_{1}$.
+The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext multiplications to compute, which is independent of the entry size, while the inner products still take $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications.
+Therefore, this optimization can reduce the computation overhead when the entry size ($t$ times the plaintext size) is larger than the first dimension size $d_{1}$.
 Keyword PIR.
@@
 We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [8] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
-We add a conditional, lazy modular reduce to Harvey’s NTT [28].
-The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size.
+We add a conditional, lazy modular reduction to Harvey’s NTT [25].
+The main idea is that we only reduce during the butterfly if the integers would exceed the machine word size.
 This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes.

### v5

@@
 Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations.
-We use basic composition in DP [36, Lemma 2.3].
+We use basic composition in DP [35, Lemma 2.3].
 Lemma 1.
@@
 Private information retrieval.
-Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation.
-Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database.
-PIR protocols can be for index retrieval or retrieval by keyword [12].
-In this paper, we focus on single-server PIR protocols [30] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard.
-We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [5] and [4].
-There is a notion of differentially-private PIR in the multi-server setting [34, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy.
-This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
+Though the paper is mostly presented for semantic search, Wally applies to keyword search as well by using keyword variant of private information retrieval (PIR).
+PIR protocols are cryptographic client-server protocols where a server holds a large database and the client can privately retrieves an entry corresponding to a keyword from the database.
+In this paper, we focus on protocols based on a single server [29].
+There is a also notion of differentially-private PIR in the multi-server setting [33, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy.
+We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
 3 System Overview
@@
 Appendix A Optimizations
-Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR).
+Here we present our optimizations for both use cases: private nearest neighbor search and exact retrieval (PIR).
 Our optimizations revolve around the BFV [9, 17] and BGV [10] SHE schemes.
@@
 We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$.
-Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
-A.1 Approximate Nearest Neighbors Optimizations
-The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size.
+Concretely, we are able to drop around $8$kB from the response size in private nearest neighbor search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
+A.1 Private Nearest Neighbor Search Optimizations
+The homomorphic computation in the private nearest neighbor search use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size.
 We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers.
@@
 The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$.
-Kim et al [29, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
+Kim et al [28, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
 Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.

### v6

@@
 Abstract
-This paper presents Wally, a private search system that supports efficient semantic and keyword search queries against large databases.
-When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems.
-In previous private search systems, for each client query, the server must perform at least one expensive cryptographic operation per database entry.
-As a result, performance degraded proportionally with the number of entries in the database.
-In Wally, we remove this limitation.
-Specifically, for each query the server performs cryptographic operations against only a few database entries.
-We achieve these results by requiring each client to add a few fake queries, and send each query via an anonymous network to the server at independently chosen random instants.
-Additionally, each client also uses somewhat homomorphic encryption (SHE) to hide whether a query is real or fake.
-Wally provides $(\epsilon,\delta)$-differential privacy guarantee, which is an accepted standard for strong privacy.
+This paper presents Wally, a private search system that supports efficient search queries against large databases.
+When sufficiently many clients are making queries, Wally’s performance is significantly better than previous systems while providing a standard privacy guarantee of $(\epsilon,\delta)$-differential privacy.
+Specifically, for a database with 3.2 million entries, Wally’s queries per second (QPS) is 7-28x higher, and communication is 6.69-31x smaller than Tiptoe, a state-of-the-art private search system.
+In Wally, each client adds a few fake queries and sends each query via an anonymous network to the server at independently chosen random instants.
+We also use somewhat homomorphic encryption (SHE) to reduce the communication size.
 The number of fake queries each client makes depends inversely on the number of clients making queries.
-Therefore, the fake queries’ overhead vanishes as the number of clients increases, enabling scalability to millions of queries and large databases.
-Concretely, Wally can process eight million queries in 117 minutes, or just under two hours.
-That is around four orders of magnitude faster than the state of the art.
-1 Introduction
-Consider a scenario where a client holding a search query wants to retrieve relevant information from a large database hosted on a server.
+Therefore, the overhead of fake queries vanishes as the number of honest clients increases, enabling scalability to millions of queries and large databases.
+I Introduction
+Consider a scenario where a client holding a potentially private search query wants to search information from a large database hosted on a server.
 This scenario reflects the flow of current search engines such as Google, Bing, and DuckDuckGo.
@@
 Therefore, enabling query privacy is critical for search engines.
-Private search engines opt to provide query privacy.
-However, these systems have significantly higher computational and communication overhead than the state-of-the-art insecure solutions.
-For example, in Tiptoe [25], a state-of-the-art private search engine, a single client query is 21 MB and requires 339 core-seconds of server computing against a database of $400$ million entries.
+Private search systems prioritize user privacy.
+There is a long list of works that have explored protecting query privacy, by periodically injecting fake queries that resemble real queries [1, 2, 3, 4, 5] or obfuscating users’ original queries using decoy terms [6, 7, 8, 9].
+These approaches are quite efficient.
+However, they lack a standard privacy guarantee and often rely on assumptions that limit adversarial capabilities.
+Many systems assume adversaries have restricted knowledge—whether unaware of the obfuscation algorithm, lacking background data, or using a predefined inference strategy.
+As a result, when these assumptions are removed or a more capable adversary is considered, fake queries become identifiable, ultimately breaking query privacy [10, 11, 12, 13].
+Recent privacy search systems do provide the strongest privacy guarantee - the server does not learn any information about the client query irrespective of knowledge and capability, but these schemes incur very high server computation and communication which makes them impractical.
+For example, Tiptoe [14], a state-of-the-art private search system, requires a communication of around 17 MB and provides queries per second (QPS) of only 909, with 10,000 cores, for a database with 3.2 million entries.
 The high computational cost is because, for each client query, the server must scan the entire database; otherwise, it will learn which database entries the client is not interested in.
-Similarly, the client must send cryptographic material for each database entry, which results in high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick.
-We detail this trade off in a later section..
-An additional limitation of Tiptoe is that it requires each client to store a database-dependent state of around 50 MB for each client query.
-In this paper, we present Wally, a privacy preserving search engine that efficiently scales to the performance constraints of large-scale search engines.
-Wally supports semantic search, where the client wants to fetch the most relevant documents to a query, and keyword search, where the client wants to retrieve documents associated with a private keyword.
-To provide privacy, our key observation is that in search engines, there are always enough clients making queries.
-We exploit this observation to gain privacy by devising a way to hide a particular query in the crowd of queries.
-In our protocol, the server computation and the communication are amortized over the number of clients making the requests; therefore, when the number of clients is large, the per-query cost becomes minimal.
-Wally achieves these results by relaxing its privacy notion to $(\epsilon,\delta)$-differential privacy.
-In Wally, the server learns a noised differentially private histogram of anonymized queries from all clients, over a partition of the server database.
-We emphasize that previous works, e.g., Tiptoe, are fully oblivious.
-That is, the server learns nothing besides that a particular client is making a query.
-Therefore, the privacy guarantee of schemes like Tiptoe is stronger than Wally.
+Similarly, the client must send cryptographic material for each database entry, which results in a high communication cost [^1] In reality, these systems trade off response size with request size by using a standard database clustering trick.
+We detail this trade-off in a later section..
+The world of private search thus is in an undesirable state.
+On one hand, traditional approaches do not provide a standard privacy guarantee, making them vulnerable; on the other hand, schemes that provide a strong privacy guarantee have very high computational and communication requirements, due to which they cannot be used in large-scale systems.
+Our system.
+We present Wally, a privacy preserving search system that provides standard privacy guarantees while maintaining efficiency.
+Our key observation is that in search engines, there are always enough clients making queries.
+Using this observation we explore a new trade-off between efficiency and privacy.
+Specifically, in Wally some small information is leaked to the server, but this leakage is proven to preserve differential privacy for the participating users.
+In return, Wally’s computational and communication overhead reduces significantly compared to the state of the art.
+Specifically, for a database with 3.2 million entries, Wally’s QPS is 7$\sim$28x higher and communication is 6.69$\sim$31x smaller than Tiptoe.
+Additionally, the search quality (presented as MRR@@@@100 scores) of Wally is decently better than Tiptoe.
+Wally relaxes its privacy notion to $(\epsilon,\delta)$-differential privacy.
+The privacy guarantee of schemes like Tiptoe is stronger than Wally.
 However, differential privacy is an accepted standard for strong privacy [15].
-Specifically, every client in Wally is guaranteed that any statistical inferences are insensitive to their data.
-In other words, their participation does not significantly alter the server’s view, except with probability $\delta$, and thus provides strong privacy protection.
-At a high level, Wally works in epochs.
-Within each epoch, a fixed number of users make queries.
-The epoch length is picked to ensure enough clients are present to make queries, and their queries will be processed by the end of an epoch.
-The server divides the database into $K$ clusters.
-Each query consists of a relevant cluster index, and the server response is all the entries in the cluster.
-The queries arrive at the server via an anonymous network, ensuring all identifiable information like IP address is removed from the queries.
+Intuitively, every client in Wally is guaranteed that leakage to the server does not change by too much when their input is modified or removed.
+More formally, the server’s ability to infer their queries does not improve by more than a factor of $e^{\epsilon}$ , except with probability at most $\delta$.
+We emphasize that we do not make any other assumption about the client queries or the adversary’s capabilities.
+Additionally, the privacy guarantee only gracefully degrades when the adversary sees query history.
+Summary of techniques.
+Like Tiptoe, Wally also uses standard search techniques based on semantic embeddings.
+At a high level, the embedding model maps any text string into a floating-point vector of dimension $d$ with a semantic guarantee: embeddings corresponding to similar texts will have a high inner product score (cosine similarity).
+This way, the search problem is reduced to the problem of private nearest neighbors: The client must find the embedding vector that maximizes the inner product with the query vector.
+The server divides the database into $K$ clusters and sends cluster centroids to each client.
+The system works in epochs.
+The epoch length is picked to ensure that enough honest clients are present to make queries and there are enough server resources to process all the queries by the end of an epoch.
+Within each epoch, a large number of honest users make queries.
+We discuss this assumption in more detail towards the end of the introduction.
+For each query, the client finds $\Delta$ closest clusters and forwards them to the server.
+The server responds by sending all the embeddings in those clusters.
+The queries arrive at the server via an anonymous network, ensuring that all identifiable information, such as IP address, is removed from them.
 The server then generates a response to each query independently and sends it back to the respective client via an anonymous network.
-Even though queries are anonymized, they do not provide query privacy.
-The server can infer information about the client query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries).
-To prevent this leakage, Wally implements the following two changes at the client.
-First, to ensure that the server does not learn information by observing the access patterns, each client in Wally makes fake queries.
-For each fake query, the client picks a randomly selected bucket.
-This ensures that the server cannot differentiate between real and fake queries.
-The number of fake queries is sampled from a particular distribution to ensure the revealed histogram of queries is $(\epsilon,\delta)$-differentially private.
-In expectation, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries where $U$ is the number of clients within an epoch, and $\Delta$ is the maximum number of queries each client can make in an epoch.
-Note that increasing the number of clients decreases the average number of fake queries.
-The overhead due to fake queries can be traded with privacy because more fake queries are required as $\epsilon$ decreases.
+Although queries are anonymized, this does not guarantee query privacy.
+The server can infer information about the client’s query by exploiting the access pattern (if a particular cluster is accessed) or arrival time (when the server receives queries).
+To prevent this, Wally implements the following two changes at the client.
+First, each client in Wally makes fake queries.
+To generate a single fake query, the client picks a random cluster.
+The client does not have to interact with the server to generate random queries and generates them independently.
+The number of fake queries is sampled from a negative binomial distribution to ensure $(\epsilon,\delta)$-differential privacy.
+Looking ahead, to maintain privacy, the client must send only a single cluster per query.
 We emphasize that fake queries do not affect the accuracy or correctness seen by any client, as the client can discard the responses due to these queries.
-Second, Wally also ensures that arrival time and order of the queries leak no information about the client to the server.
-To achieve this, instead of making a query instantly, the clients make queries at random one second slots within an epoch.
-Also, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
-Anonymizing the queries, adding fake queries that are indistinguishable from real ones, and randomizing query order and time guarantees differential privacy.
+Second, Wally also ensures that neither the arrival time nor the order of the queries leak any information about the client to the server.
+To achieve this, instead of sending a query instantly, each client sends queries in randomly sampled one-second slots within an epoch.
+Additionally, an anonymous network collects the requests received within a particular slot and sends them as a batch to the server.
+We show that anonymizing the queries, adding fake queries, and randomizing the query schedule is equivalent to sharing a noised histogram of queries from all the clients with the server.
+We then prove that this histogram guarantees $(\epsilon,\delta)$ privacy.
 As mentioned above, for each query, the server response contains all the entries in the cluster.
-For large clusters, this would result in very high response overhead.
-In Wally, we utilize lattice-based, somewhat homomorphic encryption (SHE) to reduce the response overhead.
-Specifically, for semantic search, each query additionally includes a semantic information (dependent on the client’s input) encrypted under SHE.
-The server computes the distance function between the encrypted information and the cluster entries under SHE and returns encrypted scores.
-This reduces the response because the size of encrypted scores is significantly smaller than the entries.
-For a keyword search, the client sends an encrypted PIR query, and the server runs an SHE-based keyword private information retrieval (PIR) protocol within the cluster, which returns data associated with the client keyword.
-To ensure that the server cannot differentiate between real and fake queries, for every fake query the client includes encrypted semantic, or keyword queries based on fake data.
-We propose various optimizations that significantly improve performance for both kinds of searches.
-Overall, for a database with 35 million entries Wally can process eight million real queries in 117 minutes on 320 cores, a drastic improvement over the state-of-the-art, which would have required years to process the same queries.
-Additionally, the size of each request is 53 times smaller and the size of each response is 3.5 times smaller compared to Tiptoe.
-Our contributions are as follows:
-  • We describe Wally’s design, which enables privacy-preserving large-scale search engines.
-Its novelty stems from the combination of anonymity, fake queries added by the clients, and the client’s use of SHE to encrypt its queries.
-  • We also describe the implementation of Wally, showing efficiency and scalability along with a robust privacy guarantee.
-Generally, SHE computations have a high performance overhead.
-To achieve desired performance constraints, we propose several optimizations to SHE computation (distance computation and PIR) that could be of independent interest.
-  • We evaluate Wally for private image search, demonstrating both strong privacy and low overhead.
-Limitation.
-Wally requires a set number of clients always present in an epoch to make queries.
+For large clusters, this would result in a high response overhead.
+We utilize lattice-based, homomorphic encryption (SHE) to reduce the response overhead.
+Specifically, for each real query, the client freshly encrypts the query embedding under SHE and sends it along with the corresponding cluster to the server.
+For a fake query, the client encrypts just a zero embedding.
+To compute the distance score, the server computes the dot-product between the encrypted embeddings and the cluster embeddings under SHE.
+This reduces the response because the size of encrypted distance scores is significantly smaller than the embedding size.
+Formal contributions. Our contributions are as follows:
+  • An efficient private search system, named Wally.
+A novel system that meticulously achieves $(\epsilon,\delta)$ differential privacy for client queries.
+It achieves this by carefully stitching various steps to enhance privacy.
+Moreover, we provides a formal security proof that demonstrates the protocol’s security against a semi-honest adversary.
+  • Novel optimizations enable efficient computation of distance scores under SHE.
+  • A standalone open-source Swift library that includes an implementation of the BFV SHE scheme and various privacy primitives, including Private dot-product and Keyword Private Information Retrieval.
+This library is highly versatile and can be easily integrated into various privacy applications.
+  • We compare Wally with two state-of-art schemes and demonstrate how Wally provides a better trade-off between privacy and performance.
+Epoch size in Wally.
+Wally works in epochs, and we assume there are enough clients present in an epoch.
+Note that large epochs allow considering more clients.
 This assumption aligns with current search engines.
-For example, hundreds of millions of people make queries in Google search daily [18].
-We remark that for systems with few users, the overhead due to fake queries could get quite high, therefore fully oblivious protocols, like Tiptoe, for these systems would be more suitable.
-2 Background
+For example, hundreds of millions of people make queries in current search systems daily [16].
+So we could consider a day-size epoch with millions of users.
+However, note that epoch size also dictates the client experience; long epochs mean some clients will have to wait till the end of epoch to get response to their queries.
+Picking the very short epoch is not desirable as well.
+As few clients will present, this means high overhead due to fake queries.
+Concretely, within an epoch, on average, each client makes $O(\frac{C\Delta\log(1/\delta)}{U\epsilon})$ fake queries, where $U$ is the number of clients within an epoch, and $C$ is the number of clusters on the server.
+Moreover, $\Delta$ is the maximum number of clusters each client can probe in an epoch.
+Our experiments assume epochs of length 1-10 minutes, with that at least 100,000 to 500,000 clients are present to make queries within any epoch.
+Malicious clients can affect privacy by not submitting the fake queries.
+In our privacy calculation, we assume that at least half the clients are honest and exclude fake queries from malicious clients.
+We remark that Tiptoe does not have this requirement.
+Therefore, for systems with few clients (like a few hundred users), the overhead from fake queries will be relatively high; therefore, for such a situation, Tiptoe would be more suitable.
+We further discuss these assumptions in Section VIII.
+II Background
 Let $[n]:=\{1,\dots,n\}$ be the set of the first $n$ natural numbers starting from one.
@@
 All logarithms are the natural log unless noted otherwise.
-We denote a random variable being sampled from a probability distribution $D$ as $a\leftarrow D$.
+We denote a random variable $a$ being sampled from a probability distribution $D$ as $a\leftarrow D$.
 All distributions in this work are over finite sets and $a\leftarrow S$ denotes uniformly sampling from $S$ when $S$ is a discrete set.
@@
 Definition 1.
-A distribution $\mathcal{D}$ is infinitely-divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d.
+A distribution $\mathcal{D}$ is infinitely divisible if for all $n\in\mathbb{N}$, it can be expressed as the sum of $n$ i.i.d.
 variables.
 That is, there exists a distribution $\mathcal{D}_{n}$ such that $\mathcal{X}_{1}+\cdots+\mathcal{X}_{n}\sim\mathcal{D}$ and $\mathcal{X}_{i}\sim\mathcal{D}_{n}$.
-2.1 (Somewhat) Homomorphic Encryption
+II-A (Somewhat) Homomorphic Encryption
 Fully homomorphic encryption (FHE) is a special cryptosystem allowing arbitrary computation over ciphertexts.
@@
 We can homomorphically swap $(\mathbf{v},\mathbf{\bar{v}})$ with the same operation, called “conjugation”.
-In Table 1 we show concrete run times and noise growth for BFV, an efficient SHE scheme.
+In Table I we show concrete run times and noise growth for BFV, an efficient SHE scheme.
 In general, addition and plaintext multiplication are efficient and addition has minimal noise growth.
@@
 Evaluation keys.
-Ciphertext rotation and multiplication require evaluation keys that depend on a secret key.
-Additionally, each rotation requires a separate evaluation key.
-Usually, these keys can be reused, but as mentioned later, in Wally  the client must send fresh keys for each query.
+BFV multiplication, rotation, and substitution operations require evaluation keys that depend on the client’s secret key.
+The server could store the evaluation keys once in typical SHE applications and reuse them across various client requests.
+However, the server learns the requests belonging to the same client.
+Looking ahead, this will break the privacy guarantee in Wally  therefore, the client must send fresh evaluation keys with each request.
 Therefore, these keys contribute to the client’s request size.
@@
 $\mathsf{CtRotate}$ $0.5$ $0.5$
-Table 1: Experimental computation cost and noise growth of each BFV homomorphic operation.
+TABLE I: Experimental computation cost and noise growth of each BFV homomorphic operation.
 The polynomial degree $n$ is $4096$, the ciphertext modulus $Q$ has $83$ bits, and the plaintext modulus $t=40961$ (16 bits).
@@
 Time costs are measured with the swift-homomorphic-encryption library [19] (git commit b70d927) on an Intel Xeon w3-2423 using a single core.
+0
 Differential privacy.
-We use the standard notion of differential privacy (DP) [14, 13]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private if for all $\epsilon\geq 0$ and $\delta\in[0,1]$, any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
+We use the standard notion of differential privacy (DP) [20, 21]: a randomized mechanism $M$ is $(\epsilon,\delta)$-differentially private for $\epsilon\geq 0$ and $\delta\in[0,1]$, if for any two datasets $X,X^{\prime}$ differing on one element, and for any subset $S$ of possible transcripts output by $M$, $\Pr[M(X)\in S]\leq e^{\epsilon}\cdot\Pr[M(X^{\prime})\in S]+\delta$ where the probability is over the randomness of $M$.
 Let $\zeta$ be an output of a mechanism $M$, then the privacy loss in observing $\zeta$ between two datasets $X,X^{\prime}$ as input is
@@
 Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations.
-We use basic composition in DP [35, Lemma 2.3].
+We use basic composition in DP [22, Lemma 2.3].
 Lemma 1.
@@
 Anonymous networks (ANs), also called anonymization networks and anonymous communication protocols, are protocols which offer user anonymity by using cryptographic techniques such as onion routing [23], or mixnets [24].
-Throughout the paper, we assume that ANs 1) strip all identifying information from incoming messages from clients, and 2) permute all received messages within a batch of messages from clients before sending them to the server.
+Throughout the paper, we assume that ANs 1) strip all identifying information from incoming queries from clients, and 2) collect all of them received in a second and send them in a batch to the server.
 Private information retrieval.
@@
 In this paper, we focus on protocols based on a single server [25].
-There is a also notion of differentially-private PIR in the multi-server setting [33, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy.
+There is also a notion of differentially-private PIR in the multi-server setting [26, 27], where the database is stored across non-colluding servers while guaranteeing differential privacy.
 We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
-3 System Overview
-Figure 1: High-level overview of Wally.
-Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
-Figure 1 shows a high-level design of Wally.
-It consists of three entities: the server, the clients, and the anonymization network.
-Wally runs in epochs of time; within each epoch, a fixed number of clients make queries.
-Within each epoch, Wally performs a search that outputs the most relevant results to client queries, similar to insecure search engines.
-Wally does not require any synchronization across clients.
-Each client can independently decide to participate or skip an epoch.
-However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch.
-Overall, the privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated.
-Wally’s privacy notion also requires that each client makes at most $M$ queries in the system’s lifetime; therefore, a client can participate at most $M$ epochs.
-The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
-In Wally each epoch is divided into time slots, each a second long, e.g., and privacy is guaranteed by the following:
-  • Hiding the content of the queries (somewhat homomorphic encryption).
-  • Hiding the real queries of clients among fake queries (differential privacy).
-  • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
-  • Hiding the information revealed by queries order and time of arrival (random query schedule).
-Embedding-based search.
-Similar to state of the art insecure semantic search and previous secure search, Wally utilizes (vector) embeddings for semantic search.
-Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats.
+II-B Threat Model
+We break Wally’s entities into three groups: clients, the anonymous network’s (AN) hops, and the server.
+The server and AN hops are modeled as semi-honest adversaries attempting to learn a client’s query.
+Clients can behave maliciously, intentionally deviating from the protocol, in an attempt to learn an honest client’s query.
+The server’s database is assumed to be public.
+Assumptions.
+  • Clients: We assume there are at least two honest clients.
+Note that malicious client could break privacy of the protocol by looking ahead, as described in Theorem 1 and in VIII.
+Wally needs many more honest clients for better performance, as looking ahead since Wally relies on differential privacy.
+Malicious clients can collude with each other, an anonymous network hop, or the server.
+If they collude with an anonymous network hop, then we assume they collude with the server and a hop that is not the first hop.
+  • Anonymous network hops: We assume there are at least two hops, and that at least one hop is honest.
+This is a standard assumption in anonymous networks.
+For example, oblivious HTTP [28] has two hops where the second hop is assumed to be operated by the server and the first hop is assumed to be operated independently by a different party.
+  • Server: The server can collude with any anonymous network hop besides the entry hop, as well as clients.
+We assume the entry hop cannot collude with the server since the server would learn which client sent which queries, fake and real.
+  • Cryptographic assumptions: Lastly, we assume the CPA security of RLWE-based SHE [29, 17].
+Client privacy.
+Under the above assumptions, Wally protects client query utilizing two privacy notions:
+  • The client’s queries in Wally are protected under CPA-security of Wally’s SHE scheme.
+  • Which clusters the client queries is protected under a differential privacy guarantee, Theorem 1.
+Overall, Wally provides a differential privacy guarantee on the client’s queries because of the latter bullet above.
+Further, we make no assumptions on the server’s knowledge of the expected query distribution across all clients.
+III Private search with inefficient communication
+Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “removes” identifying information.
+In this section we provide a warm-up protocol that provides the required privacy but has high communication overhead.
+This protocol act as stepping stone towards our final protocol discussed in Section V.
+Embedding.
+Similar to state of the art insecure semantic search and previous secure search, This protocol utilizes embeddings for search.
+Embeddings are a machine learning-based technique that map unstructured objects like text, images, and videos to $d$-dimensional vectors of floating-point numbers.
 These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space.
 Hence, they are used to find relevant entries in search systems.
-Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication.
 Embeddings can be generated for different data types, including text, image, and video.
-Therefore, Wally can support all these data types.
-Wally is compatible with all the recent embedding models, such as BERT [12].
-In the following sections, we introduce the insecure version of semantic search in Wally.
-We then highlight the privacy issues in this version.
-Finally, we discuss how we overcome these privacy issues.
-3.1 Insecure Search Pipeline
+Therefore, our protocol can support all these data types.
+Wally is compatible with recent embedding models, such as BERT [30].
+Clustering based nearest neighbors search.
+The server has a database of $N$ entries, where each entry consists of a document such as text or image and a bit string of associated metadata.
 In the offline phase, the server maps each document to the corresponding embedding.
-At the time of a query the client embeds the query into an embedding.
-Once the client query and the server documents are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query.
-Our two-step approximate nearest neighbor protocol has the following flow:
-  • Server initialization.
-At initialization, the server divides the documents (using their embeddings) into $K$ clusters using K-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$.
+At the time of a query, the client embeds the query into an embedding.
+Once the client’s query and the server’s documents are represented as embeddings, the client’s objective is to identify the nearest server embedding and its associated metadata.
+Finding the exact nearest neighbor is computationally expensive even in an insecure environment.
+Therefore, we utilize an approximate nearest neighbor algorithm using clustering to find the most relevant documents to the client’s query.
+The protocol has the following flow:
+  • Initialization.
+At initialization, the server divides the document embeddings into $K$ clusters using $K$-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$.
+The server also divides the document metadata into the same clusters.
 The server sends the centroid embeddings to the client.
-  • Pick nearest centroid.
-To make a query, the client first locally finds the cluster nearest to the query embedding.
-For this, the client computes a similarity between the query and each of the centroid embeddings and picks a cluster whose centroid has high similarity.
-  • Server computation.
-The client then sends the nearest cluster index $C^{\prime}$ and query embedding $q^{\prime}$ to the server.
-The server computes similarity between the query embedding and each embedding in cluster $C^{\prime}$.
-The server returns the similarity scores to the client.
-  • Pick nearest entry. The client sorts these scores and picks the entry closest to the query.
-  • Metadata retrieval. The client then fetches the metadata associated with the closest entry from the server.
-The clustering technique provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size).
-However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is in a different cluster than the cluster with the closest centroid.
+  • Query.
+To make a query, the client first locally finds $\Delta$ clusters nearest to the query embedding $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$.
+For this, the client computes a cosine similarity between the query and each of the centroid embeddings and picks clusters whose centroid has the highest similarity.
+  • Response.
+The client then sends the nearest clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the server.
+The server responds by sending the embeddings and their metadata in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the client.
+  • Local computation.
+The client locally computes cosine similarity between the query embedding and each embedding in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{1}={\Delta}$, sorts the scores to find the closest embedding, and the corresponding metadata.
+The clustering provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size).
+However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is missed.
 Therefore, finding a balance is important for the overall usability of the system.
-The above protocol is not private. To ensure that the protocol does not reveal any information about the client’s query, it must hide the following information from the server:
-[…]
-  • Nearest cluster.
-The client reveals to the server the nearest cluster ID.
-Recall that clustering puts highly correlated documents together in a cluster.
-Therefore, the cluster ID can reveal the kind of data the client holds.
-For example, if the client query is a song embedding, the cluster ID might reveal the genre of the song.
-  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
-We show how we hide the above data below and in Section 6.
-3.2 Hiding the Query Embedding
-We rely on efficient vectorized, RLWE-based SHE described in Section 2 to hide a query embedding.
-At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster.
-The server then returns encrypted scores to the client.
-SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient.
-Overall, in our implementation, the client query and server response are a single RLWE ciphertext of sizes 226kB, and 53kB, respectively, (the former with evaluation keys) and it only takes 19 ms for the server to compute similarity scores for a single cluster homomorphically.
-3.3 Hiding the nearest cluster
-By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster.
-Hiding the nearest cluster is a significant performance bottleneck in previous schemes like Tiptoe.
-Specifically, to conceal the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database (all the clusters).
-Otherwise, skipping clusters will reveal to the server that the client is not interested in them.
-Computing over the entire database hides the nearest cluster in the most robust possible sense, but this results in a prohibitively high server computation cost.
-Instead, we propose an approach based on differential privacy.
-The key advantage is that, in expectation, the server computes encrypted similarity scores only against a few clusters per query.
-Our approach results in a significant improvement in server computation.
-An additional advantage of our approach is a reduction in request size; concretely, the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe.
-Conversely, differential privacy is a weaker privacy notion than the full, cryptographic obliviousness Tiptoe achieves.
-The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making the queries at any given time.
-Therefore, we propose a novel mechanism to hide a particular client query among the crowd of queries from all the clients.
-Our proposed approach provides $(\epsilon,\delta)$-differential privacy.
-This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $\epsilon$ with all but $\delta$ probability.
+This protocol is not private.
+To prevent the protocol from disclosing any information about the client’s query, it must conceal the closest clusters from the server.
+Remember that clustering groups highly correlated documents together in a cluster.
+Consequently, knowing the nearest clusters could reveal the information about the client’s query.
+III-A Hiding nearest clusters using differential privacy
+Hiding the nearest clusters is a significant performance bottleneck in previous schemes like Tiptoe.
+For instance, to conceal a single nearest cluster, Tiptoe requires the client to send an encrypted query for each cluster; similarly, it requires the server to perform SHE operations against all the clusters.
+This is because omitting clusters will indicate to the server that the client is not interested in them.
+Computing over the entire database hides the nearest cluster completely, but this results in a prohibitively high communication and computation.
+We propose an approach based on differential privacy.
+The key advantage is that touching all the clusters is not required, which results in a significant improvement in communication and server computation than Tiptoe.
+We emphasize that differential privacy is a different privacy notion than the full obliviousness Tiptoe achieves; however, it is still a standard privacy notion.
+Intuitively, our proposed approach provides $(\epsilon,\delta)$-differential privacy.
+This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $e^{\epsilon}$ with all but $\delta$ probability.
 In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
-As the first step, we require all the traffic (queries and responses) to and from the server to route through the anonymization network.
-Intuitively, one would expect that doing this would hide the query in the crowd of queries, as queries do not carry identifiable information.
-However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can learn information about a client.
-For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
+The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making queries at any given time.
+Therefore, we can hide a particular client query among the batch of queries from all the clients.
+Our approach require following steps.
+Epochs.
+As the first, we require the system to run in epochs of time, each one second long; within each epoch, a fixed number of clients make queries.
+This does not require any synchronization across clients.
+Each client can independently decide to participate or skip an epoch.
+However, we assume a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch.
+We also require that the client send $\Delta$ clusters as independent queries to the server instead of sending them together.
+The worst privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated.
+The correctness guarantee is that all the queries within an epoch are processed by the end of the epoch.
+Anonymized queries.
+We also require all the queries to be anonymized. For this, each client will send the queries via an anonymization network.
+Note that even though queries are anonymized, by observing traffic patterns, the server can still learn information about the queries.
+For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not.
+Therefore, we require deniability of queries.
 Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client.
 For example, the queries towards each epoch’s start could belong to a particular user.
-To achieve the desired privacy, we make the following two additional changes, highlighted in Figure 3:
+Therefore, we need to suppress leakage through the timing channel.
+To achieve the desired properties, we make the following two additional changes, highlighted in Figure 3:
   • Fake queries.
@@
 Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated.
-In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries.
-For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks one of the clusters.
+In expectation, the client makes $K\leftarrow\mathbb{E}[\mathcal{D}]$ fake queries..
   • Random query schedule.
@@
 Intuitively, these two changes suppress the leakage, as mentioned above.
-In Section 6, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
-3.4 Hiding Metadata Index
-After learning the index of the nearest entry, the client retrieves metadata associated with it.
-When metadata per entry is small, the server returns metadata of the entire cluster with each query.
-Note that this leaks no extra information because the server already learns the cluster ID for each query.
-The client then locally selects the metadata associated with the most relevant entry.
-However, if the metadata is quite large, the client can run Keyword PIR to fetch the metadata of only the most relevant entry within a cluster.
-3.5 Keyword Search
-Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword.
-Using a random hash function, the server divides the database into $K^{\prime}$ random buckets (instead of clusters).
-For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket.
-The differential privacy part remains the same as in semantic search.
-That is, clients make fake queries for random buckets at a random schedule.
-4 Efficient SHE-Based Instantiation
-In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls.
-In this section, we discuss the details of our efficient SHE-based instantiation.
-4.1 SHE details in semantic search
-The server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
+In Section VI, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees $(\epsilon,\delta)$ differential privacy.
+Drawback of warm-up protocol.
+Even though the protocol conceals the closest clusters, it incurs high communication costs.
+Concretely, the server transmits embeddings and metadata for all the nearest clusters back to the client, leading to substantial communication overhead.
+IV Optimizing the communication using SHE
+We use SHE to reduce the communication of our warm-up protocol. In this section, we discuss the details of our efficient SHE-based instantiation.
+IV-A Reducing communication in nearest neighbor
+[…]
+We rely on vectorized, RLWE-based SHE described in Section II to hide a query embedding.
+At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster.
+The server then returns encrypted scores to the client.
+SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient.
 Server database.
@@
 Figure 2: A single slice computation in Wally using SHE operations. The server only require aligning the query once for all the slices.
-SHE dot-product computation.
-Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them.
+Secure dot-product computation.
+From Section II, to calculate cosine similarity between two normalized embeddings, the server needs to compute a dot-product between them.
 To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$.
@@
 Again, for the databases we consider, $\lceil N^{\prime}/n\rceil\approx 1$ results in $d$ multiplications, and each response is a single ciphertext.
-Recall from Section 2 that homomorphic rotation is costly.
+Recall from Section II that homomorphic rotation is costly.
 In the above explanation, the server performs $d$ rotations even after packing to align the query.
@@
 See Algorithm 1 in [32] for the BSGS algorithm.
-Float-to-integer encoding.
-Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats.
-In Wally we convert each float into an integer by multiplying it with the appropriate scaling factor having the plaintext modulus be large enough to perform the computation over the integers.
-The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
 Figure 3: Differential privacy related changes to hide real queries. Changes are highlighted in gray.
-4.2 SHE details in keyword search
-The server database consists of $N$ (keyword, value) entries.
-Recall that the server divides these entries into $B$ buckets using a random hash function.
-Within each bucket, in Wally  we use a keyword PIR construction based on cuckoo hashing and index PIR.
-At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions.
-The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table.
-The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data.
+IV-B Metadata Fetch
+After determining the index of the nearest entry, the client retrieves metadata related to it.
+When the metadata per entry is small, the server responds by returning the metadata for the entire cluster with each query.
+It’s important to note that this doesn’t leak any additional information since the server already knows the query cluster.
+The client then locally selects the metadata associated with the most relevant entry.
+However, if the metadata is large (in kilobytes), it becomes infeasible to download the entire cluster.
+In this case, the client obliviously retrieves the metadata of only the most relevant entry within a single cluster.
+To achieve this, we use Keyword private information retrieval (PIR) scheme based on cuckoo hashing and index PIR.
+Concretely, after getting encrypted scores, the client finds the most relevant embedding locally.
+Use the identity of this entry to generate a Keyword PIR query.
+The metadata query is then clustered in which this embedding falls, as well as the Keyword PIR query.
+Protecting privacy leakage.
+Note that, it is still required to protect the query cluster.
+Therefore, clients will require participating in an additional epoch with $\Delta=1$.
+For this epoch also, we required queries to be anonymized, clients generate fake queries, and execute queries at random slots within an epoch.
+Details of Keyword PIR.
+The server encode each entry into Keyword-value pair.
+Where keyword equals entry ID and value is the metadata.
+For ease of explanation we assume that each cluster on average consist of $N^{\prime}$ entries.
+The server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions.
+The guarantee is that if a keyword is present in the database, it must be at entry index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table.
+The client then uses index PIR to privately fetch entries at these two indices and locally find the index containing the entry and associated data.
 Wally uses MulPIR [33] as the underlying index PIR, an efficient single-server PIR scheme based on SHE.
 We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
-Each server bucket consisting of $N/B$ entries is represented as a $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$.
-MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial.
-For this explanation, we assume that each entry has the same size as SHE plaintext.
+Details of MulPIR.
+The server represent cuckoo table into $\sqrt{N^{\prime}}\times\sqrt{N^{\prime}}$ matrix For this explanation, we assume that each metadata entry has the same size as SHE plaintext.
 If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable.
@@
   1. 1.
-    First expand $\hat{q}$ to two $\sqrt{N/B}$-length ciphertext vectors.
+    First expand $\hat{q}$ to two $\sqrt{N^{\prime}}$-length ciphertext vectors.
 Encrypting row indicator vector $\mathbf{r}$ and column indicator vector $\mathbf{c}$.
 The server uses the oblivious expansion algorithm given in [33] to expand.
-This algorithm requires $2\sqrt{N/B}$ homomorphic substitutions.
+This algorithm requires $\sqrt{N^{\prime}}$ homomorphic substitutions.
   2. 2.
-    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N/B$ plaintext-ciphertext multiplications and $N/B-\sqrt{N/B}$ ciphertext additions.
+    Then compute the matrix-vector product $\mathbf{a}:=\mathbf{D}\mathbf{c}$ with $N$ plaintext-ciphertext multiplications and ciphertext additions.
   3. 3.
-    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response.
-This step uses $\sqrt{N/B}$ ciphertext-ciphertext multiplications and $\sqrt{N/B}-1$ ciphertext additions.
-to compute.
+    Finish by computing $\mathbf{r}^{T}\mathbf{a}$ that results a single ciphertext that is a response. This step uses $\sqrt{N^{\prime}}$ ciphertext-ciphertext multiplications and additions.
 Section A discusses various optimizations we implemented for MulPIR. Overall, our optimizations resulted in a three-fold reduction in server computation and a three-fold reduction in request size.
@@
 18:Output $D_{\text{sem}},c$
-5 Wally Protocol
+V Wally Protocol
 In this section, we describe the complete Wally protocol for a single epoch in semantic search.
@@
 7:return $r:=(R,\textsf{data})$
-6 Security
+VI Security
 In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch.
@@
 Second, we show that the view of the server in an epoch can be simulated only using the curator’s output.
-6.1 DP Security in the Central Model
+VI-A DP Security in the Central Model
 We define the central model as follows:
@@
 We emphasize that the theorem is stated for a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
-6.2 Generating view of the server
+VI-B Generating view of the server
 We show how the curator’s output is sufficient to simulate the server’s view generated by distributed DP algorithm defined in Algorithm 2 together with an anonymization network (AN).
@@
 Therefore, the query schedule generated in both Hybrids is indistinguishable.
-To complete the proof, note that in Line 3, the simulator generates $\mathsf{list}$ in the same way as the curator in the central model.
-After that, the simulator uses this $\mathsf{list}$ to generate a random query schedule, which results in a server’s view indistinguishable from the view in Hybrid 0.
-7 Implementation and Performance Evaluation
-In this section, we demonstrate Wally’s concrete performance via a set of experiments.
-We also showcase Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
-7.1 Implementation
-We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [6].
-Following the library, we use BFV [17] to implement the SHE parts of Wally.
-However, our protocol could easily be implemented using other SHE schemes such as BGV [10].
-BFV parameters
-The choice of BFV parameters provides a trade-off between computation and communication in Wally.
-At a high level, larger parameters provide better computation but result in high request size.
-We tried multiple sets of parameters and picked the one that provided the best performance.
-We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits.
-These parameters allow a precision of six bits.
-To increase the precision, we use plaintext CRT optimization described in Section A.
-This optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query).
-For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$.
-This gives us a precision of 13 bits.
-Differential privacy parameters
-Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries.
-For a fixed failure probability $\delta$, smaller $\epsilon$ provides stronger privacy but worse performance because each client is required to make more fake queries.
-We set $\delta=2^{-20}$, which means DP will fail only once a million and $\epsilon=1$, providing a decent privacy guarantee.
-7.2 Performance Evaluation
+To complete the proof, observe that in Line 3 of Hybrid 2, the simulator generates $\mathsf{list}$ in a manner similar to the curator in the central model.
+However, the key difference lies in the scheduling of queries.
+The simulator sends queries from $\mathsf{list}$ at random intervals, while the curator in the central model sends $\mathsf{list}$ all at once to the server.
+Despite the adversary potentially recognizing its interaction with the simulator, it gains no additional advantage because the random schedule remains independent of secret inputs from the user.
+VII Implementation
+We have implemented Wally in Swift and Python languages. We next discuss implementation details for each component of Wally
+VII-A Open Source SHE library
+Components of Wally outlined in Section IV are instantiated using the Brakerski-Fan-Vercauteren (BFV) encryption scheme [17].
+We have released the SHE components of Wally as an open-source Swift library named Apple Swift Homomorphic Encryption Library available at https://github.com/apple/swift-homomorphic-encryption.
+It includes the BFV scheme, a secure dot-product for search, and Keyword PIR for metadata fetch.
+We have implemented many optimizations within the library, which significantly improve the overall performance and can be of independent interest.
+The library API is simple and flexible, and can be used for other applications.
+We also provide an easy to follow tutorial and examples to use the API.
+We discuss a few key optimizations and implementation decisions:
+  • Residue Number System (RNS) variant of BFV.
+In BFV scheme, we manipulate elements modulo ciphertext modulus $Q$, which is an integer with hundreds of bits.
+Implementing this scheme requires slow multi-precision modulo arithmetic.
+To address this, we instead implemented the RNS variant of BFV [35].
+In this variant, $Q$ is chosen as $Q=q_{1},\ldots,q_{l}$, and each $q_{i}$ is a single-precision integer of typically 28 or 55 bits.
+By using the RNS, any integer $x\in Z_{Q}$ can be represented as $\{x_{1}=x\mod q_{1},\cdots,x_{l}=x\mod q_{l}\}$ and the operation on $x\mod Q$ can be performed by performing the same operation on each RNS component $x_{i}$ modulus $q_{i}$.
+At the end we can use CRT to extract original ciphertext in $Z_{Q}$.
+We also implemented the Hybrid RNS variant of key-switching [36] and fast base conversion [37] methods required for $\mathsf{CtCtMult}$ and $\mathsf{CtRotate}$ operations.
+  • Reducing the cost of BFV operations.
+Number Theoretic Transformation (NTT) speeds up polynomial multiplications in BFV operations.
+We implemented NTT using Harvey’s butterfly [38] and optimized it using lazy modular reduction.
+In every butterfly operation, we track multiplications and only perform modulus when the result overflows the machine word size.
+This optimization improves forward and backward NTT by $20-25\%$.
+In BFV multiplication, we perform tensor product and addition in large basis, then scale it back to modulus $Q$.
+When performing many BFV multiplications, we skip the scaling step after each multiplication and perform it only once at the end [39].
+This results in $10-20\%$ improvement in MulPIR used for fetching metadata.
+  • Reducing the size of response ciphertexts.
+At the end of the Wally protocol, the server sends response ciphertexts back to the client.
+The client just decrypts these ciphertexts and will not use them for further computation.
+Due to this, we can use the standard modulus switching technique to reduce the size of these ciphertexts by scaling down to the RNS limb $q_{l}$.
+In BFV after mod switch, the message is encoded into higher $\log_{2}{t}$ bits.
+Therefore, we can drop the least significant $\log_{2}{q_{l}/t}$ bits of the response ciphertexts and still correctly decrypt the message.
+However, doing so does add some noise in the resulting ciphertext.
+We carefully monitor this noise and drop the bits that still allow correct decryption.
+In [40], the authors suggested that we can drop $l_{0}$ and $l_{1}$ bits from the first and second ciphertext polynomials as long as $z\sqrt{2\frac{n}{9}}2^{l_{1}}+2^{l_{0}}<2^{\log_{2}{(q_{l}/t)}}$, where $z$ represents a score of uniform distribution.
+The decryption error is bounded with z score probability.
+We picked $z=8$ that gives decryption error of $2^{-37.5}$.
+We set $l_{0}=\log_{2}{(q_{l}/t)}-2$, and calculates $l_{1}\leq\lfloor\log_{2}(q/t)-2-\log_{2}(z\sqrt{2n/9})\rfloor$.
+If $l_{1}\leq 1$ we update $l_{0}=l-1$ and $l_{1}=0$.
+    Note that we assume that the original error in the ciphertext is smaller than the error introduced by dropping bits.
+  • Reducing the request size.
+As mentioned in Section II, evaluation keys are essential for SHE operations.
+In more detail, during the search, the server utilizes evaluation keys to rotate and merge ciphertexts after multiplication.
+To minimize key size, the client transmits only two keys in two rotations: one for the giant step and another for the baby step (the server repeatedly uses the same keys for all other rotations).
+Furthermore, we introduce an RNS-based plaintext decomposition method that maintains a constant key size while achieving high precision.
+Specifically, we select the modulus $t^{\prime}$ (which allows high precision without wrap-around; more details in the next section) as a multiple of two smaller NTT-friendly primes, $t_{0}$ and $t_{1}$.
+The server decomposes each database vector into two RNS components.
+The client decomposes the query into two RNS components, encrypts each separately using the same secret key $s$ but different plaintext modulus as $t_{1}$.
+The server performs exact dot-product computations using each query ciphertext, and the server returns both of them to the client.
+After decryption, the client reconstructs the original plaintext in $Z_{t}$ using the Chinese Remainder Theorem (CRT).
+Our main observation is that the evaluation keys do not depend on the plaintext modulus used in the ciphertext; therefore, the server will use the same evaluation key for both queries.
+    In metadata fetch, to expand a query, the server performs repeated substitution operations on the query ciphertext.
+Typically, the server requires one substitution key for each required power.
+However, we propose a method where the client sends keys for only a few substitutions, and the server computes the remaining substitutions using the same keys.
+We provide detailed explanations in the Appendix.
+VII-B Embedding and Clustering
+We map each document in the server database into dense embeddings that retain semantic information.
+For this, we use the sentence embedding model ‘msmarco-distilbert-base-tas-b’ with 66 million parameters and optimized for semantic search [^3] https://huggingface.co/sentence-transformers/msmarco-distilbert-base-tas-b.
+In particular, we use the implementation provided by the ‘sentence-transformers’ Python library [^4] https://www.sbert.net.
+The model outputs dense 768-dimensional floating point vectors.
+We then use PCA to reduce the dimensionality of each vector to 192.
+The next step is clustering these vectors.
+To do so, we first find $K$ cluster centroids by running k-mean clustering.
+Then we assign each embedding the cluster with the nearest centroid.
+Finally, all points and their assigned cluster are compiled into an index.
+We implemented this pipeline using the open-source library FAISS.
+FAISS provides fast, GPU-based algorithms for k-means clustering and indexing.
+For a comprehensive list of optimizations implemented by FAISS, refer to https://github.com/facebookresearch/faiss/wiki#research-foundations-of-faiss.
+FAISS exposes a flexible clustering API that allows us to experiment with various clustering parameters, such as the number of data points sampled, the number of iterations of the basic k-means algorithm, and whether to run on GPU or CPU.
+For indexing, we utilize one of the Flat indexes provided by FAISS to perform fast intra-cluster nearest neighbor search.
+VII-C Parameters Selection
+SHE parameters.
+Performance of Wally directly depends on our choice of BFV parameters.
+We tried multiple parameter sets and selected the one that yielded the optimal performance.
+For search and metadata fetch, the selected parameters are as follows: polynomial degree $n=4096$, ciphertext modulus of $\log_{2}{Q}=83$ bits, with CRT limbs of $27,28$  and $28$ bits.
+We sample error from a centered binomial distribution with standard deviation $3.2$ and secret key from a ternary distribution modulus $Q$.
+This achieves 128 bits of quantum security.
+For search, we set plaintext modulus of $\log_{2}{t}=16$ bits.
+Recall that the embedding model outputs floating-point vectors of dimension $d$, however BFV operates over integer vectors.
+We convert each float into a fixed-point integer by multiplying it by the appropriate scaling factor $p$ and then rounding down to the nearest integer.
+We select $p$ to ensure the dot products do not wrap around modulo $t$.
+We say that precision is equal to $\log_{2}(p)$ bits.
+To avoid wrap-around, it is important that $p<\sqrt{(t-1)/2}-\sqrt{d}/2$, which implies that $\log_{2}(p)<1/2\log_{2}t\approx 7$.
+Therefore, we obtain at least 7 bits of precision.
+To increase the precision, we use the plaintext CRT decomposition.
+For this, we pick two plaintext CRT limbs of plaintext modulus of $\log_{2}{t_{0}}=16\text{ and }\log_{2}{t_{1}}=17$ bits.
+Total plaintext bits are 31 bits, which gives us a precision of 15 bits without increasing evaluation key size.
+As mentioned above, we reduce the size of the response ciphertext by first mod switching down to the last CRT limb of 27 bits and then further dropping the least significant bits.
+To avoid decryption error due to dropping bits, we set $l_{0}=9$ and $l_{1}=0$.
+This drops the size of the single response ciphertext from 55 KB to 23.5 KB.
+For MulPir to handle larger noise growth, we use a plaintext modulus of $\log_{2}{t}=5$ bits.
+Having a smaller plaintext modulus allows us to set $l_{0}=19$ and $l_{1}=13$, which reduces each response ciphertext to 12 KB.
+Epoch and number of users.
+As described in Section III, Wally operates in discrete epochs.
+In each epoch, a total of $U$ users submit queries.
+However, malicious users can undermine Wally’s privacy guarantees by refraining from submitting fake queries.
+For our privacy analysis, we assume that half of the total clients in any epoch behave maliciously.
+We compute the privacy loss by excluding fake queries from these malicious clients.
+In other words we consider worst-case behavior of all the malicious clients.
+Longer epochs lead to a larger $U$, which in turn reduces the number of fake queries required per user.
+However, extending the epoch duration may negatively impact service quality, since some queries might only be processed at the epoch’s end.
+For our evaluation, we use a epoch size of one and ten minutes and assume that at least 100,000 and 500,$0$ users submit queries during each epoch, of which half are honest.
+Given that search engines like Google and Bing process millions of queries per minute and serve billions of users, 500k is a conservative estimate in the context of modern search engines [41].
+We also assume that each user will participate in around 400 epochs, one or two epochs each day of a year.
+Differential privacy parameters.
+For a fixed failure probability $\delta$, a smaller $\epsilon$ offers stronger privacy but compromises performance since each client must execute more fake queries.
+We set $\delta=2^{-30}$, ensuring that DP will fail only once in a billion queries, and $\epsilon=1/800$, providing a reasonable trade-off between privacy and performance.
+In each epoch, we assume that each participating client makes $\Delta$ queries.
+These queries represent the number of nearest clusters fetched for each client’s query.
+As mentioned in Section VI, this guarantees $(1/400,2^{-26})$ differential privacy per epoch.
+Consequently, the total differential privacy for each client across 400 epochs is $(1,2^{-26})$.
+For each cluster, a client samples fake queries from $\mathcal{NB}(r/U,p)$. Which means in expectation the number of fake queries is
+$\mathbb{E}[\text{Fake queries per client}]=\frac{rpK}{(1-p)U}$ (1)
+$=\frac{3(1+\log_{2}(1/2^{-30}))e^{-0.0005/\Delta}K}{1-e^{-0.0005/\Delta}U}$ (2)
+where $K$ is the total number of clusters and $U$ is the total number of participating honest clients.
+A little bit work shows that the above equation is always $\leq 102,700\Delta K/U$.
+As mentioned above considering honest users $U=250,000$, a modest estimate given modern search engines’ scale [41].
+Therefore number of fake queries is upper bounded by $\leq 0.205\Delta K$
+Search parameters and metric
+The product of clusters $K$ and the number of clusters probed $\Delta$ determine the performance and accuracy of Wally.
+A smaller $K$ and a smaller $\Delta$ both enhance performance by reducing the overhead of fake queries.
+Further notice that very small $K$ will result in large response size.
+Similarly, reducing $\Delta$ may lead to a substantial drop in accuracy due to the increased probability of missing the true nearest neighbor.
+Our objective is to minimize the fake queries overhead, which is proportional to the product $K\Delta$, while balancing search quality and response size.
+However, to achieve this using exhaustive search over a candidate set of parameters would require considerable time.
+We experimented with different parameters and selected those that provide best accuracy while keeping $K$ moderately small and minimizing $\Delta$ as much as possible.
+For each probe, the client generates a new query with a large evaluation key.
+Consequently, large $\Delta$ will result in large total request size.
+Dataset
+We use the MSMARCO document ranking dataset  [42].
+The dataset contains approximately 3.2 million passages, where each passage comprises a document identifier, document URL, document title, and document body.
+VIII Evaluation
+In this section, we demonstrate Wally’s concrete performance and compare it with baselines.
 Experimental setup.
 We ran server-side computations on an Intel Xeon w3-2423 instance with 32GB of RAM and 6 cores.
-End-to-end performance.
-Metric Low prec. Mid prec. Tiptoe
-Query Size (MB) 0.67 0.84 16.2
-Response Size
-(MB) 0.16 0.18 3.2
-Time to process
-queries 117 mins 139 mins 2.8 years
-Precision (bits) 6 13 4
-Privacy DP DP SHE
-Table 2: Performance comparison of Wally and Tiptoe for database of 35 millions entries.
-Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query).
-Throughput for Wally is measured for 8 million real queries, or 24 million in total.
-All Wally responses are computed as the encrypted scores plus $44$ KB of metadata.
-Table 2 describes our main experiments compared to Tiptoe.
-For an end-to-end system, we estimate the network time on a 30Mbps channel.
-This comes out to 74ms for transmitting a query and response over the network for the small precision setting (first column in Tables 2 and 3).
-The server database consists of 35 million entries divided into around 8500 clusters.
-We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$.
-We assume each client makes $\Delta=2$ queries per epoch for these experiments.
-Moreover, there are four million honest clients present in an epoch.
-This results in eight million real queries in an epoch.
-For Wally, we calculated the overhead by using Theorem 2. The fake query overhead is derived by the expected number of fake queries for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
-$\mathbb{E}[\text{total fake queries per client}]=\frac{rp}{1-p}\cdot\frac{K}{U}$ (1)
-where $K$ is the total number of clusters and $U$ is the total number of clients.
-For our selected parameters, for each real query, the client is required to make just under two fake queries.
-Therefore, the total number of queries within an epoch is 24 million.
-In Table 2, we included fake query overhead in query request, response and total computation plus estimated network time in Wally.
-The epoch size in Wally is the total time it takes to process all the queries (real and fake).
-For Tiptoe, we assume there are eight million queries the server processes one by one.
-Tiptoe requires running experiments on real compute nodes; due to resource limitations, we carefully estimate their results using Table 7 in [25].
-Before discussing Wally’s performance, we emphasize that Wally requires the following two extra assumptions not required in Tiptoe.
-  • Wally requires sufficiently many honest clients during an epoch.
-  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
-In other words, Wally trades off privacy guarantees for scalability.
-Wally has significantly better performance than Tiptoe across all the metrics.
-Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe.
-This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters.
-Similarly, the response size is 3.1-3.6 smaller.
-Wally’s efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe.
-For example, to process eight million queries, the server with 320 cores requires around 117 minutes in Wally.
-In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
-Scaling DP with $\epsilon$.
-Figure 4 shows how the expected fake queries per cluster for each client scales with $\epsilon$ for one million clients and four million clients.
-This is derived by dividing Equation 1 by the number of clusters $K$.
-Note that the total number of clients has a linear effect while the dependence on $\epsilon$ is nonlinear as we decrease $\epsilon$.
-Figure 4: The relationship between the number of queries per cluster per client for $500$k and $1$M total clients. We fix the number of real queries to $\Delta=2$.
-Precision.
-Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section A.
-The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$.
-Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
-Precision (bits) 6 13 29
-Query (KB) 226 283 901
-Response (KB) 53 63 196
-Server comp. (ms) 19 35 110
-Table 3: Here we list the query and response size per each query (real and fake).
-Each response size includes $44$kB of metadata.
-Multiply all numbers by $\phi+1$ if DP necessitates $\phi$ fake queries per real query.
-Scaling with the number of clients.
-Table 4 shows how the fake query overhead varies with the number of messages per epoch $\Delta$ and the number of clients in the system.
-We emphasize that the number of fake queries increases linearly with the number of buckets and decreases proportionally to the number of clients.
-As mentioned above, this can be offset by increasing the size of the clusters to a multiple of $n$ and we can perform the similarity computation in parallel with more threads.
-Table 4 emphasizes the importance of having many clients in a system since the overhead vanishes as the number of clients becomes large.
-Number of Clients .5M 1M 2M 10M
-Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
-Exp. MB ($\Delta=1$) $1.9$ $1.08$ $0.65$ $3.11$
-Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
-Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$ $0.61$
-Table 4: This table shows how the fake query overhead changes with the number of clients and the number of messages. We fix the number of clusters to $K=8545$ to match our 35M dataset, Table 2.
-7.3 Keyword Query Performance
-Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [6].
-Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4].
-Each response size was 27KB.
-Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
-Database entries Entry Size (Bytes) Query size (KB) Time (ms)
-100k 10 368 166
-100k 100 539 465
-1 million 10 539 565
-1 million 100 624 1700
-Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section A.
-8 Related Work
-Anonymization networks alone do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
-Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness.
-Tiptoe [25], the most recent and performant, has been discussed above in-detail.
-Other less efficient solutions include Coeus [1] as well as Hers [16], with the latter providing privacy for the server’s database as well.
-Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[7, 31], and including a client public key in each query [19].
-Another related area of work is differentially-private multi-server PIR schemes [33, 2].
-In both works, servers learn differentially private, noised histograms over queries.
-The former, [33], has clients generate fake queries whereas the latter, [2], has the servers inject fake queries.
-Both require a secure multiparty computation over the entire database.
+Metrics
+To compare the performance, we use following metrics
+  • Queries per second (QPS): this metric reflects how many queries the server can process in a second for fixed infrastructure.
+We assume that the server has at least 10,000 cores.
+We measure the latency for a single query on a single core and divide the total number of cores by that number.
+  • Bandwidth: We calculate the size of the client query and the server response for each query.
+  • Accuracy: To evaluate the accuracy of our search results, we used the Mean Reciprocal Rank (MRR@100) metric, a standard measure for assessing the quality of search results.
+Given a query $q$ and a ground truth index $i$ such that $DB[i]$ is the most relevant entry, we merge all scores (along with entry indices) obtained from probing $\Delta$ clusters, sort them, and retain the first 100 scores.
+If the list contains $i$ at the j-th rank (where $j\leq 100$), the reciprocal rank score is calculated as $1/j$.
+In the absence of $i$ in the list, MRR@100 is set to 0.
+Baselines.
+We compared Wally with following baselines. Both baselines offer full obliviousness but have high performance overhead:
+  • Tiptoe (simulated) [14], which is a state of the art private search system based on linear homomorphic encryption (LWE) .
+Similar to Wally, Tiptoe uses a clustering approach to find the nearest candidate.
+However for each query they only probe a single cluster.
+Unfortunately, due to resource constraints we are unable to run the complete Tiptoe system.
+We therefore simulate it by using following estimates
+    • Latency: For latency we estimate the minimum time the server will take to process a single query.
+We exclude network cost, which gives advantage to Tiptoe as their network bandwidth is quite high.
+For each query, the computation is dominated by dot-products between vectors.
+The server first generates a token by multiplying RLWE encryption of secret key with precomputed LWE ciphertext.
+For this $16\sqrt{N}$ dot products are performed between vectors of dimension $n=2048$.
+Note that even though this computation can be performed in offline phase, the server still has to perform it for each query.
+Then to privately fetch and rank candidates from a particular cluster, the server performs $N$ dot-products between vectors of dimension $d$.
+For each vector we assume that each element is of 32 bits.
+    • Accuracy: To simulate their accuracy results we divide the database into $\sqrt{N}$ clusters, probe a nearest cluster and then calculate MRR@100 of the resulting candidates.
+    • Communication: To estimate communication we calculate minimum number of 32 bit elements the client and server have to transfer per query.
+To generate a token, the client sends at least $n=2048$ RLWE ciphertexts or $n^{2}$ elements.
+For ranking, the client sends a query consisting of $d\sqrt{N}$ elements.
+For tokens, the server response consists of $16$ RLWE ciphertexts (each consisting of at least 2048 elements) and for ranking the response consists of $\sqrt{N}$ elements.
+  • Pacmann, [43] which is a concurrent work.
+Like Wally, this scheme also employs sublinear server computation, but it offers full obliviousness.
+Pacmann’s approach differs from Wallyin two key aspects.
+Firstly, to locate the nearest neighbor, Pacmann utilize a graph-based search algorithm.
+Secondly, Pacmann employs sublinear Offline/Online PIR to retrieve graph nodes and traverse the graph.
+    For each query, the scheme necessitates multiple round trips and per-query maintenance.
+We compare our results with the ones provided in the Pacmann paper [43], which is sufficient since we conducted our own experiments in the matching environment.
+For Wally, we calculate the QPS and bandwidth by summing the cost of $\Delta$ real queries and fake queries.
+Comparison with Tiptoe.
+The results are presented in Table II.
+Wally significantly outperforms Tiptoe across all metrics.
+Notably, Wally’s QPS is 11~27 times better, this is because for each query the server has to perform at least one expensive operation per entry in the database.
+Similarly, the communication per query is 10~34 times smaller than Tiptoe, for varying numbers of clusters probed.
+Tiptoe can only probe a single cluster per query, and to probe more clusters, a new query must be initiated.
+For $\Delta=1$, a single probe case, Wally’s MRR100 score is only slightly better than Tiptoe’s score of 0.11.
+However, as we increase $\Delta=5$, Wally’s MRR100 score improves to 0.18, which is considerably better than Tiptoe’s score.
+Comparison with Pacmann.
+As given in the table, Pacmann’s scheme has better QPS and MRR100 than Wally.
+Specifically, the better QPS is because Pacmann is based on offline/online sublinear PIR and their MRR100 is better due to use of graph based search data structure which is superior than clustering technique.
+However, their scheme requires large client storage and communication per query.
+Specifically, for 3.5 million database, the client is required to store a hint of 614 MB and per query communication is 61.6 MB, which is prohibitively high, specifically for cellular networks.
+Another downside is that hint-based protocol can not directly handle database updates.
+Finally, to gain better MRR100 the scheme requires multiple round trips which adds to their network delay.
+We did not consider the cost of network delay in their QPS calculation.
+Specifically, due to their use of offline/online PIR, even though QPS of Pacmann is 1.35.6×1.3$\sim$ 5.6$\times$1.3 ∼ 5.6 × better than Wally but the communication is 23123×23$\sim$ 123$\times$23 ∼ 123 × as large for varying $\Delta$.
+Client Storage (MB) Communication (MB) Queries Per Second (QPS) MRR100
+Tiptoe [14] 0.61 17.4 909 0.11
+Pacmann [43] 614 61.6 34,482 0.26
+Wally ($K=256$, $\Delta=1$) 0.04 0.56 25,974 0.12
+Wally ($K=256$, $\Delta=3$) 0.04 1.7 9,881 0.16
+Wally ($K=256$, $\Delta=5$) 0.04 2.6 6,667 0.18
+TABLE II: Comparison of Wally  Tiptoe and Pacmann for MSMARCO [42] database.
+Communication and QPS for Wally includes overhead due to fake queries.
+Worst value in each metric is marked with a red color.
+Benchmarking Wally.
+In Table III, we evaluate the performance of Wally across various database sizes ranging from one million entries to hundred million entries and cluster sizes ranging from 128 to 1,024.
+For all these experiments, we assume that the client probes the nearest cluster $\Delta=1$.
+Additionally, we consider epoch sizes of one and ten minutes, with 100,000 and 500,000 thousand honest users, respectively.
+Firstly, we notice that increasing the epoch size (i.e., the number of honest users) leads to improved overall performance across all database sizes.
+Specifically, the QPS increases by 2-3 times, while the request and response sizes improve by 1.5-2.4 times.
+This demonstrates that Wally is an ideal choice for applications where moderate to extended delays are acceptable.
+Secondly, we observe that the expected number of fake queries is nearly proportional to the product of the number of honest users ($U$) and inverse of the number of clusters ($K$).
+For each database, as the epoch size increases from one minute to ten minutes, we increase the number of users by 5×5$\times$5 × (which reduces fake queries) but also increase the number of clusters by 2×2$\times$2 × (which increases fake queries).
+Consequently, we see that fake queries almost decrease by a factor of $5/2=2.5$, indicating a linear dependency.
+Alternatively, we can consider that reducing the number of clusters $K$ for a fixed database will lead to a decrease in fake queries.
+Thirdly, we observe that as the database size varies, the request and response sizes linearly increase.
+However, the QPS decreases.
+This is because for larger databases, each cluster size becomes substantial, resulting in increased server operations per query.
+Entries in database N 1 Million 16 Million 100 Million
+Epoch size 1 Minute 10 minutes 1 Minute 10 minutes 1 Minute 10 minutes
+Number of users $U$ 100,000 500,000 100,000 500,000 100,000 500,000
+Number of clusters (K) 128 256 256 512 512 1,024
+Expected fake queries 1.7 0.3 3.3 1.7 6.7 2.6
+Request size (MB) 0.76 0.36 1.18 0.76 2.17 1.01
+Response size (MB) 0.17 0.04 2.3 1.5 13.9 3.03
+Queries per second (in thousand) 222 666 14 32 4.7 13.9
+TABLE III: Performance of Wally across varying database sizes. Each client will probe a single cluster $\Delta=1$. Each setting provides $(\epsilon=1,\delta=2^{-26})$-DP guarantee.
+Number of clusters (K) 64 128 256 512
+Expected fake queries 0.3 0.6 1.2 2.4
+Request size (MB) 64 128 256 512
+Response size (MB) 64 128 256 512
+QPS (in thousand) 64 128 256 512
+TABLE IV: Performance of Wally for different values of $K$.
+Each epoch is of ten minutes and database has one million entries.
+Number clusters probed $\Delta=3$.
+Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
+Clusters probed ($\Delta$) 1 3 5 10
+Expected fake queries 0.3 0.6 0.8 1.3
+Request size (MB) 0.36 0.45 0.59 0.65
+Response size (MB) 0.08 0.10 0.11 0.14
+QPS (in thousand) 370 312 277 217
+TABLE V: Performance of Wally for different values of $\Delta$.
+Each epoch is of ten minutes and database has one million entries.
+Number of clusters $K=125$.
+Each setting provides $\epsilon=1,\delta=2^{-26}$ DP guarantee.
+IX Related Work
+Early private search systems.
+Early private search systems heuristically add dummy queries or obfuscate queries by changing the query client-side, but they ultimately send queries in the clear to the server [44, 45, 46, 47, 6, 48, 49, 50, 51].
+These systems are broken by a semi-honest server, i.e., the server alone can de-anonymize queries [52, 53, 54, 55] or they have strict distribution requirements on queries, e.g., i.i.d.
+queries [47].
+This contrasts with Wally  which has a provable differential privacy guarantee: the server’s view between the histograms over clusters with and without a client’s queries is differentially private.
+Further, the client’s queries in Wally are encrypted under homomorphic encryption to prevent the trivial attack shown in [26, Section 3].
+Homomorphic encryption.
+Homomorphic encryption alone is a popular technique which has been used in private search systems [14, 56, 57, 58].
+A major downside of these systems is that they require encrypted computation over the entire database, whereas Wally computes on a few clusters per query.
+Wally achieves this by utilizing differential privacy to hide the requested cluster index from the server, which makes Wally practically deployable in settings with many honest clients.
+On the other hand, these previous systems based only on homomorphic encryption have the strongest possible privacy notion: fully obliviousness from the server’s point of view assuming standard cryptographic hardness assumptions.
+An in-depth comparison with Tiptoe [14] in Section I.
+A related system is presented in [6] which sends text queries in the clear but uses additively homomorphic encryption [59] to compute an encrypted similarity score.
+The queries are noised, or “embellished”, beforehand by replacing each word in the query with a set of unrelated words.
+Embellishing heuristics provide no provable security guarantee, unlike the differential privacy guarantee in Wally.
+Secure multiparty computation.
+Secure multi-party computation (MPC) is a common technique for private search systems.
+These systems incur large query sizes or high latency due to many communication rounds.
+SANNS [60] uses garbled circuits and homomorphic encryption for private nearest neighbors search in the semi-honest client-server model.
+As a result, they have complete cryptographic obliviousness for each query and the server’s database entries are private besides the query’s response, but each query sends gigabytes of data for databases of 1M entries or more due to their use of garbled circuits [60, Table 1].
+Preco [61] is a two-server maliciously-secure MPC private nearest neighbor search system which requires an linearly-sized computation by each server ($N$), a $\sqrt{N}$ computation by the client, and $O(N)$ communication for each query [61, Table 1].
+They achieve fully cryptographic obliviousness via distributed point functions [62] and probabilistic batch codes [63].
+Concretely, they achieve 10MB of communication and 10s latency for a database with 10M entries.
+Differential privacy.
+Differentially-private multi-server (information-theoretic MPC) PIR schemes [26, 27] are not immediately applicable to private search since PIR itself only supports exact matches, but these works’ use of DP is related to our use of DP since they use DP to lower their online computation while leaking information to the servers.
+In both works, servers learn differentially private, noised histograms of queries over database entries.
+Note, this compares to Wally where the server learns noised histograms of queries over clusters and only requires one server.
+The latter [27] is more efficient and requires a multi-server PIR over the entire database once on the database for a batch of queries.
 Notably, [27] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation.
 Their offline and online computation grows linearly with the size of the server’s database.
-Differential privacy has also been applied to multiple anonymous messaging systems [36, 34, 30].
-These are large (many server) mix-net-like protocols where the servers add in fake queries to hide user messages with a differential privacy guarantee.
-Though they are for a different application, their use of differential privacy is similar to Wally’s.
-9 Conclusion
+Wally, on the other hand, has no offline computation and is sublinear in the database size.
+Anonymization networks.
+Anonymization networks and mix-nets [64, 24] anonymize the user’s identity but send the query in the clear.
+Only applying these methods to search systems would not hide users’ private information and are broken by a semi-honest server [51].
+Trusted hardware.
+Trusted hardware is another tool used in private search systems [65, 66, 67]. However, trusted hardware search systems are only as strong as their underlying hardware [68, 69, 70, 71, 72].
+Concurrent works.
+Pacmann [43] efficiently reduces graph-based private nearest neighbors search to many PIR queries on a graph.
+Pacmann uses offline-online PIR which requires the client to stream the database beforehand, then each PIR computation is sublinear since it uses a recent sublinear PIR protocol [73].
+As a result, it is a cryptographic, fully-oblivious private search system and offers strong accuracy results from their use of graph-based nearest neighbors search.
+However, Pacmann requires each client stream the entire database offline whereas Wally requires no offline computation.
+Further, Pacmann requires many rounds per query, unlike Wally which only requires one round for small database entries and two rounds for large database entries.
+X Conclusion
 Wally demonstrates that private search systems can be scalable to systems with many users.
@@
 First, we describe the standard RNS optimization in SHE [74] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [75].
-RNS form.
 Often $Q$ is larger than machine-size words and we use $Q$ as the product of machine-sized primes $Q=q_{0}q_{1}\cdots q_{l}$.
@@
 Concretely, we are able to drop around $8$ kB from the response size in private nearest neighbor search to a response size of $47$ kB when combined with our BFV plaintext CRT optimization described in the next subsection.
-A.1 Private Nearest Neighbor Search Optimizations
+A-A Private Nearest Neighbor Search Optimizations
 The homomorphic computation in the private nearest neighbor search use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size.
@@
 One limitation of this optimization is that we run out of NTT-friendly plaintext primes for small parameter sets, like $n=4096$ or $2048$.
-A.2 PIR Optimizations
+A-B PIR Optimizations
 Uneven dimensions.
 In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication.
-For example, Table 1 shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
+For example, Table I shows ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $0.5ms$ and ciphertext multiplication is $2.5ms$.
 Therefore, we structure the database (cluster) as a rectangle $\mathbf{D}\in R_{t}^{d_{1}\times d_{2}}$ with uneven dimensions $d_{1}$ and $d_{2}$.
@@
 The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$.
-Kim et al [28, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
+Kim et al [39, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one.
 Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.